### v1

# Real Noncommutative Convexity I

Abstract.
We initiate the theory of real noncommutative convex sets, following the recent and profound complex case developed by Davidson and Kennedy. The present paper focuses on the real case of the topics from the first several sections of their profound memoir [11]. Later results will be discussed in a future paper. We give many foundational structural results for real operator systems and their associated noncommutative (nc) convex sets, and elucidate how the complexification interacts with the basic convexity theory constructions. Several new features appear in the real case, including the novel notion of the complexification of a nc convex set.
1. Introduction
Let $X$ be a compact Hausdorff space. A concrete function system is a selfadjoint unital subspace $V$ of $C(X)$, where $C(X)$ is the abelian $C^{*}$-algebra of scalar valued continuous functions on $X$. Our scalar field $\mathbb{F}$ will either be $\mathbb{R}$ or $\mathbb{C}$. An element $f\in V$ is called positive if for all $x\in X$ we have $f(x)\geq 0$. A state on $V$ is a scalar valued linear functional on $V$ which is unital, selfadjoint, and positive in the sense that it maps positive functions to positive numbers. Equivalently, these are the unital contractive functionals on $V$. The collection of states on $V$, denoted by $S(V)$, is a convex set. By the Banach-Alaoglu theorem it is also compact with the $w^{*}$-topology. Conversely, given a compact convex set $K$ the set $A(K)$ of affine scalar valued functions on $K$ is a function system inside $C(K)$. There is also an abstract characterization of function systems which will be discussed in a later section.
Kadison’s representation theorem shows that there is a duality between function systems and compact convex sets. Indeed, we have that for all function systems $V$ and compact convex sets $K$
$A(S(V))\cong V,\quad S(A(K))\cong K,$
where the isomorphism in both cases is given by evaluation. The map taking $V\mapsto S(V)$ and $K\mapsto A(K)$ is a contravariant functor, so that the category of convex sets is dually equivalent to the category of function systems. All the above works for convex sets in both real or complex vector spaces, and for real or complex function systems (see e.g. [1, 22]). However the relationship between the real and complex theory is complicated in places. For example, many convexity theoretic results about complex function spaces $V$ are proved via Re $(V)$, i.e.  in the real setting, since the direct complex variant can be messy.
In the amazing paper [11], Davidson and Kennedy establish a profound noncommutative convexity theory, in the complex case. Early in this work they exhibit a categorical equivalence similar to the one above, but for noncommutative (nc) convex sets and complex operator systems, giving a non-commutative analogue of Kadison’s duality. This is built on previous work of Webster and Winkler [29], who use complex matrix convex sets instead. Almost all the matrix and noncommutative convexity hitherto was done in the complex case (see e.g. [30, 10] for the very few existing results in the real case). Recently, the first author and Russell developed the theory of real operator systems in [6]. The present paper is a sequel to this. Given any real operator system, there is a very natural way to complexify to get a complex operator system. The process of complexification is functorial in the sense that many of the constructions done with operator systems (for instance, max/min operator systems, duals, and tensor products) usually commute with complexification. This allows the complex theory of operator systems to be applied in the real case.
Therefore it is natural to ask if there is a categorical equivalence between real compact noncommutative (nc) convex sets and real operator systems, and more generally if the theory of Davidson and Kennedy carries over to the real case. There are several motivations for this. For example, classical convexity is in some ways essentially a real theory, as inspection of foundational texts such as [1, 24] shows, or e.g. as one sees in nearly all graduate courses on convexity theory. Thus the real theory is likely to play a future role in operator algebras and mathematical physics. As mentioned in more detail in the second paragraph of [6], real structure occurs naturally and crucially in very many areas of mathematics and mathematical physics, and in several deep mathematical theories at some point a crucial advance has been made by switching to the real case (e.g. in $K$-theory and the Baum-Connes conjecture, see also for example [25]). This is sometimes because the real category is bigger and hence allows more freedom. In our case, every complex operator system (resp. nc convex set) is clearly a real operator system (resp. real nc convex set), but there are many interesting real operator systems (resp. nc convex sets) that are not complex operator systems (resp. complex nc convex sets). E.g. the selfadjoint matrices; the real nc convex set associated with these is not a nc convex set in the sense of [11]. This is a somewhat trivial example, but it illustrates the point.
In this paper we initiate the theory of real nc convex sets, and investigate the real case of Davidson and Kennedy’s theory, at least up to chapter 5 of [11]. Indeed since our paper is already lengthy, later results will be discussed in a future paper. Several new features appear in the real case. Many of these are connected to the fact that, as opposed to the classical case of convex sets and function systems, it turns out that there is a very natural way to complexify a nc convex set. Because this complexification is functorial this will give us an efficient way to generalize the theory to the real case. We give many foundational structural results for real operator systems and their associated nc convex sets. In particular we elucidate how the complexification interacts with the basic convexity theory constructions. In addition, we include some results about nc convex sets in the real and complex case that do not seem to be in the literature. The differences between the real and complex case discovered in [6] show up for us too, such as the absence of a Min and Max functor for the nc convex sets corresponding to general real operator systems such as the quaternions, and that the nc $A(K)$ need not be order isomorphic to the classical variant on $K_{1}$ (the ‘first level’ of $K$).
Turning to the structure of our paper, Section 2 gives some background on real and complex operator systems and their complexifications. In keeping with the task and nature of our paper we do however expect the reader to be reading alongside with parts of [6, 11]. Because of this we also do not need to be very pedantic or overly careful with definitions, preliminaries, or the history of the subject, which may usually be found in detail there. In Section 2.2 we define real and complex non-commutative (nc) convex sets and give basic examples such as the real non-commutative state space. Section 3 deals with how we can take the complexification of a real nc convex set. This can be done intrinsically by specifying what elements will be in the complexification, or extrinsically by taking the complex nc convex hull of a real nc convex set. These two constructions will be equivalent. Here we also prove functorial properties of complexification. For instance, if $K$ is a nc convex set and $A(K)$ are the nc affine functions on $K$, then
$A(K_{c})=A(K)_{c}.$
In Section 4, we show the real version of Davidson and Kennedy’s categorical duality. This can either be done by doing their proof in the real case, using a real version of the nc separation theorem (see 3.3), or by the functoriality of complexification from Section 3.
Section 6 begins with some facts about function systems, and then discusses how a compact convex set may be turned into a compact nc convex set. This is in duality with the way a complex operator system can be given a minimum and maximum operator system structure [23, 31]. However this is different in the real case. It is shown in [6] that a real operator system $V$ can be given a minimum or a maximum operator system structure if and only if $V$ has trivial involution. Thus any real function system can be given a minimum and maximum operator system structure as done in Section 9 of [6]. The classical compact convex sets come from real compact nc convex sets corresponding to the latter class of real operator systems. As with operator systems, the min and max structure given to a real convex set commutes with complexification. This process uses the bipolar of a nc convex set, and so we develop that in Section 5. Section 7 develops the important notion of non-commutative functions in the real case. Kennedy and Davidson do this using the theory of Takesaki and Bichteler. We may avoid the latter explicitly by e.g.  proving key theorems such as $4.3.3$ in [11] in the real case by complexification.
2. Preliminaries
2.1. Operator Systems and Operator Spaces
For general background on operator systems and spaces, and in particular on the definitions etc. in the rest of this section, we refer the reader to e.g. [21, 5, 11] and in the real case to e.g. [3, 6]. It might also be helpful to also browse some of the other existing real operator space theory e.g. [26, 27, 28, 7, 4]. Some basic real $C^{*}$- and von Neumann algebra theory may be found in [18].
We write $M_{n}(\mathbb{R})$ for the real $n\times n$ matrices, or sometimes simply $M_{n}$ when the context is clear. Similarly in the complex case. We sometimes use the quaternions $\mathbb{H}$ as an example: this is simultaneously a real operator system, a real Hilbert space, and a real C-algebra, usually thought of as a real *-subalgebra of $M_{4}(\mathbb{R})$ or $M_{2}(\mathbb{C})$. Its complexification is $M_{2}(\mathbb{C})$. The letters $H,K$ are usually reserved for real or complex Hilbert spaces. Every complex Hilbert space $H$ is a real Hilbert space, i.e. we forget the complex structure. More generally we write $X_{r}$ for a complex Banach space regarded as a real Banach space. We write $X_{\rm sa}$ for the selfadjoint elements in a *-vector space $X$. In the complex case $M_{n}(X)_{\rm sa}\cong(M_{n})_{\rm sa}\otimes X_{\rm sa}$, but this fails for real spaces. A subspace of $B(H)$ is unital if contains the identity, and a map $T$ is unital if $T(1)=1$. Our identities $1$ always have norm $1$. We write ${\mathfrak{R}}\,a$ for $\frac{1}{2}(a+a^{*})$, while for $z\in M_{n}(\mathbb{C})$ we write Re $\text{Re }z$ for $x\in M_{n}(\mathbb{R})$ where $z=x+iy$ for $y\in M_{n}(\mathbb{R})$. Finally, for a cardinal $n$ we define the isometry $u_{n}=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$ where $1_{n}$ is the $n$-dimensional identity operator.
A concrete complex (resp. real) operator system $V$ is a unital selfadjoint subspace of $B(H)$ for $H$ a complex (resp. real) Hilbert space. For $n\in\mathbb{N}$ we have the identification $M_{n}(B(H))\cong B(H^{(n)})$ where $H^{(n)}$ is the $n$-fold direct sum of $H$. From this identification, $M_{n}(V)$ inherits a norm and positive cone. The latter is the set $M_{n}(V)^{+}\coloneq\{x\in M_{n}(V):x=x^{*}\geq 0$ in $B(H^{(n)})\}$. For $n\in\mathbb{N}$ we define the amplification of a linear map $\varphi:V\to W$ by
$\varphi^{(n)}:M_{n}(V)\to M_{n}(W)$
$[x_{ij}]\mapsto[\varphi(x_{ij})]$
The natural morphisms between operator systems are unital completely positive (ucp) functions, which are linear maps $\varphi:V\to W$ that are unital and every amplification is positive (or equivalently selfadjoint and contractive). The isomorphisms (resp. embeddings) of operator systems which are used in this paper are bijective (resp. injective) ucp maps whose inverse (resp. inverse in its range) is ucp. These are called unital complete order isomorphisms (resp. unital complete order embeddings); or ucoi (resp. ucoe) for short.
Similarly a concrete operator space $E$ is a subspace of $B(H)$ with norms on $M_{n}(E)$ inherited from $B(H^{(n)})$. The natural morphisms between operator spaces are the completely bounded maps, namely the linear maps $\varphi$ between operator spaces such that the amplifications of $\varphi$ are uniformly bounded. If the uniform bound is $\leq 1$ then $\varphi$ is called a complete contraction. If the amplifications of $\varphi$ are isometries then $\varphi$ is a complete isometry. The above definitions hold for both real and complex operator systems.
There are abstract characterizations of real/complex operator spaces and operator systems. An abstract real/complex operator space is a vector space $E$ with a sequence of matrix norms $\{||\cdot||_{n}\}_{n=1}^{\infty}$ satisfying Ruan’s axioms. Operator systems on the other hand need the notion of an order unit $e$: these satisfy that for any selfadjoint elements $x$ there is a $t>0$ such that $x+te\geq 0$. We say that $e$ is archimedean if $x+\epsilon e\geq 0$ for all $\epsilon>0$ implies $x\geq 0$. For abstract complex operator systems, we begin with a complex *-vector space $E$ (a vector space with a period 2 conjugate linear map $*:E\to E$), with a matrix ordering $M_{n}(E)^{+}$ and an archimedean matrix order unit $e$. The definition is the same in the real case, with conjugate linear replaced by linear. The matrix ordering consists of cones $M_{n}(E)^{+}$ in the $n\times n$ matrices of $E$, which are selfadjoint, proper, and closed under compressions by matrices $\beta\in M_{n,m}(\mathbb{C})$. An archimedean matrix order unit is an element $e\in E$ such that $e\otimes 1_{n}$ (where $1_{n}$ is the identity of the $n\times n$ matrices) is an archimedean order unit for each $n$. These conditions define an abstract operator system. One may then prove that there is a unital complete order embedding of this space into $B(H)$ for some $H$.
A real operator system can naturally be made into a complex operator system by complexification. To do this, we start with a real abstract operator system, call it $E$, with involution *, matrix ordering $M_{n}(E)^{+}$, and Archimedean matrix order unit $e$. The complexification of $E$ is the complex vector space $E_{c}$ consisting of elements $x+iy$ for $x,y\in E$. We give this a conjugate linear involution $(x+iy)^{*}=x^{*}-iy^{*}$. The matrix ordering $M_{n}(E_{c})^{+}$ will be defined by
$M_{n}(E_{c})^{+}=\{x+iy\in M_{n}(E_{c}):c(x,y)\geq 0\}$
where
$c(x,y)=\begin{bmatrix}x&-y\\ y&x\end{bmatrix}.$
The element $e+i0$ will be an archimedean order unit. With this, the complexification becomes an abstract complex operator system. Moreover one can show that this operator system structure on the complexification is the unique one satisfying Ruan’s completely reasonable condition, namely that the map $\theta_{E}(x+iy)=x-iy$ is a ucoi (or equivalently, is completely contractive).
In part of [6] it was checked that many of the basic theorems and constructions for complex operator systems also hold for real operator systems. Very many foundational structural results for real operator systems were developed, and it was shown how the complexification interacts with the basic constructions in the subject. In certain parts of our paper we will need real operator systems $V$ with trivial involution $x^{*}=x$. It is easy to see that these coincide with the operator systems which are the selfadjoint part of complex operator systems. (Or, of real operator systems.) Thus they form an important class of real operator systems. Note however that $M_{n}(V)$ has a nontrivial involution, the transpose. Just as ‘level 1’ (that is, $M_{1}(V)$) of the complex operator systems are exactly the complex function systems [22, 23], the real function systems (or real (unital) function spaces) are exactly ‘level 1’ of the real operator systems with trivial involution [6, Section 9]. At the end of Section 4 we characterize the nc convex sets associated with the selfadjoint parts of operator systems.
2.2. Noncommutative real convex sets and affine functions
As stated in e.g. [3, 6], every positive functional on a real operator system is a multiple of a state, and every contractive unital functional is a state. The norm of a positive functional (resp. cb norm of a completely positive map is its (resp. the norm of its) value at 1. The real states $\varphi$ on a real operator system $V$ are precisely the real parts of complex states on $V_{c}$ (such as $\varphi_{c}$), or of a complex $C^{*}$-algebra generated by $V_{c}$. However, the real parts of two different such complex states may coincide on $V$. Similarly, the real matrix states $\varphi$ on $V$ are precisely the ‘real parts’ Re $\text{Re }\circ\psi$ of complex matrix states $\psi$ on $V_{c}$ (such as $\varphi_{c}$).
See [5, Section 1.3] for basics about dual operator spaces and their theory. The real case is almost identical (see e.g [3]). We say a little more about the weak topology: For $E$ a real dual operator space, we have
$M_{n}(E)\cong M_{n}(CB(E_{*},\mathbb{R}))\cong CB(E_{*},M_{n})$
So, for $[f_{st}^{ij}]\in M_{n}(CB(E_{*},\mathbb{R}))$ and $[f_{st}]\in M_{n}(CB(E_{*},\mathbb{R}))$ we have that $[f^{ij}_{st}]\to[f_{st}]$ if and only if for all $[x_{kl}]\in M_{m}(E_{*})$ we have $[f_{st}^{ij}(x_{kl})]\to[f_{ij}(x_{kl})]$ in $M_{nm}$.
For a real operator space $E$, as in [11] we define $\mathcal{M}(E)=\bigsqcup_{n}\,M_{n}(E)$ (for $n$ cardinals bounded by some cardinal $\kappa$) with $M_{n}(E)$ the matrix space of $E$ for $X\subseteq\mathcal{M}(E)$. Here $\bigsqcup_{n}$ is the disjoint union. Define $X_{n}=X\bigcap M_{n}(E)$. In the case $E=\mathbb{R}$ write $\mathcal{M}=\mathcal{M}(\mathbb{R})$. A real non-commutative convex set over $E$ is a subset $K=\bigsqcup K_{n}\subseteq\mathcal{M}(E)$ such that
  1. (1)
    $K$ is graded: $K_{n}\subseteq M_{n}(E)$ for all $n$
  2. (2)
    Closed under direct sums: $\sum\alpha_{i}x_{i}\alpha_{i}^{\mathsf{T}}\in K_{n}$ for all bounded families $\{x_{i}\in K_{n_{i}}\}$ and every family of isometries $\{\alpha\in M_{n,n_{i}}\}$ where $\sum\alpha_{i}\alpha_{i}^{\mathsf{T}}=1_{n}$.
  3. (3)
    Closed under compressions: $\beta^{\mathsf{T}}x\beta\in K_{m}$ for every $x\in K_{n}$ and isometry $\beta\in M_{n,m}$.
As in [11] we say that $K$ is closed/compact if $E$ is a dual operator space and $K_{n}$ is closed/compact in the weak topology in $M_{n}(E)$.
For $\{x_{i}\in M_{n_{i}}(E)\}$ bounded and $\{\alpha_{i}\in M_{n_{i},n}(\mathbb{R})\}$ such that $\sum\alpha_{i}^{\mathsf{T}}\alpha_{i}=1_{n}$, a nc convex combination of $x_{i}$ is defined as $\sum\alpha_{i}^{\mathsf{T}}x_{i}\alpha_{i}\in M_{n}(E)$. As in the complex case (see Proposition 2.2.8 in [11]) a subset $K\subseteq\mathcal{M}(E)$ is nc convex if and only if it is closed under nc convex combinations.
Example 2.1.
Let $a,b\in\mathbb{R}$ with $a<b$. Then for $n\in\mathbb{N}$ let $K_{n}=[a1_{n},b1_{n}]$ where
$[a1_{n},b1_{n}]=\{\alpha\in(M_{n}(\mathbb{R}))_{sa}:a1_{n}\leq\alpha\leq b1_{n}\}$
$K=\bigsqcup_{n\in\mathbb{N}}K_{n}$ is a real compact convex set over $\mathbb{R}$ called the real compact operator interval. If we replace $\mathbb{R}$ in the above definition with $\mathbb{C}$ then we get the complex compact operator interval. We have $M_{n}(\mathbb{R})\subseteq M_{n}(\mathbb{C})$ and $M_{n}(\mathbb{R})^{+}\subseteq M_{n}(\mathbb{C})^{+}$ and therefore the real operator interval is a subset of the complex operator interval. Here we see that $K_{1}$ agree in the real and complex case because $K_{1}$ is the interval $[a,b]\subseteq\mathbb{R}$.
Example 2.2.
Let $V$ be a real operator system, then the real nc state space $K=\mathrm{ncS}(V)=\bigsqcup_{n}\mathrm{UCP}(V,M_{n}(\mathbb{R}))$ is a point-weak compact nc convex set over $V^{*}$. This norms $V$. Indeed as in the complex case
$\|[v_{ij}]\|=\sup\{\|[\varphi(v_{ij})]:n\in\mathbb{N},\varphi\in\mathrm{UCP}(V,M_{n}(\mathbb{R}))\}.$
To see this quickly note that taking a ucoe $\varphi:V\to B(H)\cong M_{\kappa}$ does this in one shot. For finite dimensional subspaces $K$ of $H$ the compressions $P_{K}\varphi(\dot{)}_{|K}$ achieve in the limit the norm above, identifying $B(K)\cong M_{n}$. It also follows from the later formula $V\cong A(\mathrm{ncS}(V))$.
The following proposition is useful for extending arguments about matrix convex sets to non-commutative convex sets.
Proposition 2.3.
Suppose we have $K$ a closed nc convex set over a dual operator space $E$, a net $\{x_{i}\in K_{n_{i}}\}$, and a net of isometries $\{\alpha_{i}\in\mathcal{M}_{n,n_{i}}\}$ such that $\lim\alpha_{i}\alpha_{i}^{\mathsf{T}}=1_{n}$ and $\lim\alpha_{i}x_{i}\alpha_{i}^{\mathsf{T}}=x\in M_{n}(E)$. Then, $x\in K_{n}$.
Proof.
Same as in the complex case. See Proposition 2.2.9 in [11]
This result implies as in [11, Proposition 2.2.10] that
Proposition 2.4.
Suppose that $K$ and $L$ are closed nc convex set over a dual operator space $E$. If $K_{n}=L_{n}$ for all $n<\infty$ then $K=L$.
The natural morphisms between real nc convex sets are real nc affine functions. These will be maps $\theta:K\to L$ between real nc convex sets which are graded, respect direct sums, and equivariant with respect to isometries. That is,
  1. (1)
    $\theta(K_{n})\subseteq L_{n}$
  2. (2)
    $\theta(\sum\alpha_{i}x_{i}\alpha_{i}^{\mathsf{T}})=\sum\alpha_{i}\theta(x_{i})\alpha_{i}^{\mathsf{T}}$ for all bounded families $\{x_{i}\in K_{n_{i}}\}$ and every family of isometries $\{\alpha_{i}\in M_{n,n_{i}}\}$ where $\sum\alpha_{i}\alpha_{i}^{\mathsf{T}}=1_{n}$.
  3. (3)
    $\theta(\beta^{\mathsf{T}}x\beta)=\beta^{\mathsf{T}}\theta(x)\beta$ for every $x\in K_{n}$ and isometry $\beta\in M_{n,m}$.
$\theta$ is continuous if $\theta|_{K_{n}}:K_{n}\to M_{n}(\mathbb{R})$ is continuous for every $n$. $A(K)$ is the space of all continuous affine nc functions from $K$ into $\mathcal{M}(\mathbb{R})$.
For $K,L$ classical convex sets and for $\varphi:K\to L$ a bijective function which is affine, its inverse is easily seen to be affine. The same will be true in the non-commutative case. Indeed, $\varphi^{-1}$ is graded because for $l\in M_{n}(L)$ then $\varphi^{-1}(l)\in M_{n}(K)$, and for $y\in L_{n}$ and an isometry $\beta\in M_{n,m}$ then $\beta^{\mathsf{T}}y\beta=\beta^{\mathsf{T}}\varphi(\varphi^{-1}(y))\beta=\varphi(\beta\varphi^{-1}(y)\beta)$. Taking $\varphi^{-1}$ of the left and right hand sides gives the result. A similar proof holds to show $(2)$.
2.3. Some relations between the real and complex case
Lemma 2.5.
The function Re :()():$\text{Re }:\mathcal{M}(\mathbb{C})\to\mathcal{M}(\mathbb{R})$ taking a complex matrix $A+iB$ (where $A,B\in M_{n}(\mathbb{R}))$ to the real matrix $A$ is real affine and completely contractive. The same is true for the map Im :()():$\text{Im }:\mathcal{M}(\mathbb{C})\to\mathcal{M}(\mathbb{R})$ taking $a+ib$ to $b$.
Proof.
The map Re is well defined and graded because it sends $N\times N$ matrices over the complex numbers to $N\times N$ matrices over the reals. To show (3), let $A+iB\in M_{N}(\mathbb{C})$ and $\beta\in M_{N,M}(\mathbb{R})$ be an isometry. Then we have
Re ($\text{Re }(\beta^{\mathsf{T}}(A+iB)\beta)$ $=\text{Re }(\beta^{\mathsf{T}}A\beta+i\beta^{\mathsf{T}}B\beta)$
$=\beta^{\mathsf{T}}A\beta=\beta^{\mathsf{T}}\text{Re }(A+iB)\beta$
The same proof holds to show condition (2) and for the imaginary part. ∎
If $V=M_{n}(\mathbb{C})_{\rm sa}$ then $V_{c}$ may be identified (via a unital complex complete order isomorphism, and of course complete isometry) with a canonical subspace of $M_{n}(\mathbb{C})\oplus M_{n}(\mathbb{C})$. Namely if $x,y\in V$ then $z=x+iy$ in $V_{c}$ is identified with $(z,z^{T})\in M_{n}(\mathbb{C})\oplus M_{n}(\mathbb{C})$. For a general complex operator system $W,$ if $V=W_{\rm sa}$ then the canonical complex linear map $u:V_{c}\to W$ is an isometric and unital identification, since $x+iy$ for $x,y\in V$ in both cases may be identified with $c(x,y)\in M_{2}(W)$. We claim that $z\mapsto(u(z),\overline{u(z)^{*}})$ is a unital complete order embedding $V_{c}\to W\oplus^{\infty}\bar{W}$. Indeed since $V$ is a real subsystem of $W_{r}$, we have $V_{c}$ a complex subsystem of $W_{c}=W\oplus^{\infty}\bar{W}$. The embedding of $W_{c}$ into $W\oplus^{\infty}\bar{W}$ is $x+iy\mapsto(x+iy,\overline{x-iy})$, which on $V_{c}$ equals $(u(z),\overline{u(z)^{*}})$.
Lemma 2.6.
For a complex operator system $V$ the ‘identity map’ taking $x+iy\in(V_{\rm sa})_{c}$ to $x+iy\in V$, for $x,y\in V_{\rm sa}$, is ucp, and is a complex linear bijective isometric order isomorphism.
Proof.
The complexification of the inclusion $V_{\rm sa}\to V$ is a canonical unital completely isometric complex map $(V_{\rm sa})_{c}\to V_{c}$. If we compose this with the canonical complex quotient map $V_{c}\to V$ (see e.g. the third paragraph of [6, Section 11]), we obtain a ucp map $(V_{\rm sa})_{c}\to V$. This agrees with the ‘identity map’. It clearly is complex linear and bijective. To see that it is an isometric order isomorphism note that $x+iy$ for $x,y\in V_{\rm sa}$ in both cases may be identified with $c(x,y)\in M_{2}(V)$. ∎
Note that the above is an order isomorphism, but not necessarily a complete order isomorphism.
Lemma 2.7.
For a complex operator system $V$ the complex nc state space of $V$ is real nc affinely homeomorphic to the closed nc subset $\{\varphi\in{\rm ncS}_{\mathbb{R}}(V):\varphi(i1)=0\}$ in the real nc state space, via the ‘real part’ operation.
Proof.
[…]
$\epsilon(\varphi)(x)=\varphi(x)-i\varphi(ix)=2(\varphi_{c}\circ j)(x),\qquad\varphi\in{\rm ncS}_{\mathbb{R}}(V),$
[…]
2.4. Affine Maps as an Operator System
Let $K$ be a real compact nc convex set and let $A(K)$ be the collection of continuous real affine maps from $K$ to $\mathcal{M}(\mathbb{R})$. As in the complex case this is a *-vector space with adjoint given by $f^{*}(k)=f(k)^{\mathsf{T}}$ for $f\in A(K)$ and $k\in K$. We identify $M_{n}(A(K))$ and $A(K,M_{n}(\mathbb{R}))$. We define the positive cone $M_{n}(A(K))^{+}$ by saying $[f_{ij}]\in M_{n}(A(K))$ is positive if and only if $[f_{ij}(k)]$ is positive for all $k\in K$.
The matrix order unit will be the constant function $1$ which sends everything in $K$ to the corresponding identity in $\mathcal{M}(\mathbb{R})$. This is a matrix order unit because an element $a$ of $M_{n}(A(K))_{\rm sa}$ will be bounded by a number $c<\infty$ by the proof in Proposition 2.5.3 of [11], which is the same in the real case. Hence $-cI\leq a(k)\leq cI$ for each $k\in K$, so that $c1-a\geq 0$. Suppose that for some $n\in\mathbb{N}$ and $f\in M_{n}(A(K))_{\rm sa}$ we have $f+\epsilon 1_{n}\geq 0$ for all $\epsilon$. Evaluating at $k\in K$ we get $f(k)+\epsilon I\geq 0$. Taking $\epsilon$ to zero shows that $f(k)\geq 0$ for all $k$, and so $f$ is positive. Therefore, $1$ is an archimedian matrix order unit and $A(K)$ is a real operator system.
Remark. Because $M_{n}(\mathbb{R})\subseteq M_{n}(\mathbb{C})$, every complex nc convex set $K\subseteq\bigsqcup M_{n}(E)$ can be regarded as a real nc convex set $K\subseteq\bigsqcup M_{n}(E_{r})$, where $E_{r}$ is $E$ regarded as a real vector space. Note that in this case complex affine functions with domain $K$ are real affine. We saw that the real nc affine functions on $K$ are a real operator system, and it contains the complex affine functions as a real subsystem.
3. Complexification
Given an operator system $E$, we define its complexification $E_{c}$ to have matrix norms $M_{n}(E_{c})$ inherited from the embedding $c:M_{n}(E_{c})\to M_{2n}(E)$
$c:[x_{nm}+iy_{nm}]\mapsto\begin{bmatrix}[x_{nm}]&-[y_{nm}]\\ [y_{nm}]&[x_{nm}]\end{bmatrix}$
If $E$ is a dual space then $E_{c}$ will be too because $E_{c}=((E_{*})^{*})_{c}=((E_{*})_{c})^{*}$.
Let $K$ be a real nc convex subset of $E$. Define the complexification of K as the set $K_{c}\subseteq\bigsqcup M_{n}(E_{c})$ by $[z_{ij}]\in(K_{c})_{n}$ if and only if $c([z_{ij}])\in K_{2n}$.
Theorem 3.1.
Given a real nc convex set $K\subseteq\bigsqcup E_{n}$, the complexification $K_{c}\subseteq\bigsqcup(E_{c})_{n}$ is a complex nc convex set with $K$ canonically embedded in $K_{c}$ via a real continuous nc affine map $\iota$. We have that $K_{c}={\rm co}_{\mathbb{C}}(\iota(K))$, the noncommutative convex hull. Also, $K_{c}$ is reasonable in the sense that $x+iy\in K_{c}$ if and only if $x-iy\in K_{c}$, for $x,y\in M_{n}(E)$. Moreover, $K$ is closed if and only if $K_{c}$ is closed.
Proof.
$K_{c}$ is clearly graded. To show (2) and (3) we need the map $c$ to behave well. Specifically, if $[x_{nm}+y_{nm}]\in(K_{c})_{N}$ and $[a_{nm}+ib_{nm}]\in M_{K,N}(\mathbb{C})$ where $a_{nm},b_{nm}\in\mathbb{R}$, then
$c([a_{nm}+ib_{nm}][x_{nm}+iy_{nm}])$ $=c([\sum_{k}a_{nk}x_{km}-b_{nk}y_{km}+ib_{nk}x_{km}+ia_{nk}y_{km}])$
$=\sum_{k}\begin{bmatrix}[a_{nk}x_{km}-b_{nk}y_{km}]&-[b_{nk}x_{km}+a_{nk}y_{km}]\\ [b_{nk}x_{km}+a_{nk}y_{km}]&[a_{nk}x_{km}-b_{nk}y_{km}]\end{bmatrix}$
$=\begin{bmatrix}[a_{nm}]&-[b_{nm}]\\ [b_{nm}]&[a_{nm}]\end{bmatrix}\begin{bmatrix}[x_{nm}]&-[y_{nm}]\\ [y_{nm}]&[x_{nm}]\end{bmatrix}$
$=c([a_{nm}+ib_{nm}])c([x_{nm}+iy_{nm}]).$
We also have
$c([a_{nm}+ib_{nm}]^{*})$ $=c([a_{mn}-ib_{mn}])$
$=\begin{bmatrix}[a_{mn}]&[b_{mn}]\\ -[b_{mn}]&[a_{mn}]\end{bmatrix}$
$=\begin{bmatrix}[a_{nm}]^{\mathsf{T}}&[b_{nm}]^{\mathsf{T}}\\ -[b_{nm}]^{\mathsf{T}}&[a_{nm}]^{\mathsf{T}}\end{bmatrix}$
$=c([a_{nm}+ib_{nm}])^{\mathsf{T}}$
Let $x_{i}\in(K_{c})_{n_{i}}$ and $\alpha_{i}\in M_{n,n_{i}}(\mathbb{C})$ be a family of isometries such that $\sum\alpha_{i}\alpha_{i}^{*}=1_{n}$. Then we have $c(\sum\alpha_{i}x_{i}\alpha_{i}^{*})=\sum c(\alpha_{i})c(x_{i})c(\alpha_{i})^{\mathsf{T}}$ where $c(\alpha_{i})$ will be a family of real isometries such that $c(\alpha_{i})c(\alpha_{i})^{\mathsf{T}}$ sum to 1. So, $c(\sum\alpha_{i}x_{i}\alpha_{i}^{*})\in K_{2n}$ which means $\sum\alpha_{i}x_{i}\alpha_{i}^{*}\in(K_{c})_{n}$. This verifies condition (2). A similar proof works for condition (3) showing $K_{c}$ is a complex noncommutative convex set. Therefore, the complexification of a real nc convex set is a complex nc convex set.
[…]
Clearly ${\rm co}_{\mathbb{C}}(\iota(K))\subseteq K_{c}$ since $K_{c}$ is convex and contains $\iota(K)$. For the reverse inequality, if $x+iy\in K_{c}$ then $x=u^{*}i_{2n}(c(x,y))u$ where u is the isometry $1/\sqrt{2}[iI_{n}\;I_{n}]^{\mathsf{T}}$. This is a nc convex combination of an element from $\iota(K)$.
The assertion about reasonability is clear from the definitions, the nc convexity, and the fact that $c(x,-y)=uc(x,y)u$ where $u$ is the unitary $I\oplus(-I)$.
Finally, suppose that $E$ is a dual real operator space and each $K_{n}$ is closed in the weak* topology. Suppose that $(x_{t}+iy_{t})$ is a net in $(K_{c})_{n}$ with weak* limit $x+iy$ in $M_{n}(E_{c})\cong M_{n}(E)_{c}$. Then $x_{t}\to x$ and $y_{y}\to y$ weak* (see [3, Lemma 5.2] and its proof). So $(c(x_{t},y_{t}))$ is a net in $K_{2n}$ with weak* limit $c(x,y)$. Thus $c(x,y)\in K_{2n}$ and $x+iy\in(K_{c})_{n}$. Thus $K_{c}$ is closed. The converse is easier. (Suppose that $(x_{t})$ is a net in $K_{n}$ with weak* limit $x$. Then $(\iota(x_{t}))$ is a net in $(K_{c})_{n}$, and $\iota(x_{t})\to x$ weak* in $M_{n}(E_{c})$.) ∎
Remark. Similar considerations show that if $K_{c}$ is nc convex then so is $K$.
We may thus define $\theta_{K}:K_{c}\to K_{c}$ as the restriction of the canonical period 2 automorphism $\theta_{E}$ of $E_{c}$ taking $x+iy\to x-iy$. Then $\theta_{K}$ is easily seen to be a period 2 real affine homeomorphism of $K_{c}$ whose fixed points are $K$. Conversely if $C$ is a complex nc convex set in $E_{c}$ possessing a period 2 real affine homeomorphism, then the set $K$ of its fixed points is easily seen to be a real nc convex set.
Example 3.5 shows that complexification can be complicated, and for instance change the first layer of a nc convex set quite a bit. For now we give a simple example.
Example 3.2.
Consider the real compact operator interval from 2.1. The complexification of the real operator interval will be the complex compact operator interval. To see this, first let $a,b$ be real numbers. Let $\bigsqcup[a1_{n},b1_{n}]_{\mathbb{R}}$ be the real operator interval and $\bigsqcup[a1_{n},b1_{n}]_{\mathbb{C}}$ be the complex operator interval. Take $x+iy\in[a1_{n},b1_{n}]_{c}$ and we want to show it is in $[a1_{n},b1_{n}]_{\mathbb{C}}$. By the definition of the complexification, $c(x+iy)\in[a1_{2n},b1_{2n}]$ and so we have
$a1_{2n}\leq\begin{bmatrix}x&-y\\ y&x\end{bmatrix}\leq b1_{2n}$
However, the above will hold if and only if
$a1_{n}\leq x+iy\leq b1_{n}$
and therefore $x+iy\in[a1_{2n},b1_{2n}]_{\mathbb{C}}$. Conversely, if we take $z\in[a1_{n},b1_{n}]_{\mathbb{C}}$ then $z$ can be written as the sum of a real and imaginary part, say $z=x+iy$, satisfying the centered equation above and therefore $z\in[a1_{n},b1_{n}]_{c}$.
As in Theorem 2.4.1 of [11] we have a real noncommutative separation theorem.
Theorem 3.3.
Let $K$ be a real closed nc convex set over a real dual operator space $E$. Suppose there is an $n$ and $y\in M_{n}(E)$ such that $y\not\in K_{n}$. Then there exists $\gamma\in M_{n}(\mathbb{R})_{\rm sa}$ and a normal completely bounded map $\varphi:E\to M_{n}(\mathbb{R})$ such that ${\mathfrak{R}}\,\varphi_{n}(y)\not\leq 1_{n}\otimes\gamma$ but for all $p$ and $x\in K_{p}$ we have ${\mathfrak{R}}\,\varphi_{p}(x)\leq 1_{p}\otimes\gamma$. If $0_{E}\in K$ we can take $\gamma=I_{n}$. If $E$ is a real operator system and $K\bigcup\{y\}$ consists of selfadjoint elements, then $\varphi$ can be chosen to be selfadjoint.
Proof.
This follows as in the complex case in [11] from the Effros-Winkler separation theorem in [13, Theorem 5.4]. The real version of the latter is proved exactly as in the complex case (needing the real version of the GNS representation theorem [18, Theorem 3.3.4] applied to a faithful real state of $M_{n}$). ∎
Let $V$ be a real operator system and consider $\mathrm{ncS}(V)_{c}$. This will be a complex nc convex set, however so is $\mathrm{ncS}(V_{c})$. There is a canonical map
$\psi:\mathrm{ncS}(V)_{c}\to\mathrm{ncS}(V_{c})$
induced by the canonical isomorphism $CB(V,W)_{c}\cong CB(V_{c},W_{c})$ (see for example [4, Theorem 2.3]). Indeed for $x,y\in V$ and $f,g\in M_{n}(\mathrm{CB}(V,\mathbb{C}))$ we have
$\psi(f+ig)(x+iy)=f(x)-g(y)+if(y)+ig(x).$
The inverse map takes $u\in CB(V_{c},W_{c})$ to ${\rm Re}\,u_{|V}+i{\rm Im}\,u_{|V}$, where ${\rm Re},{\rm Im}$ here denote the two canonical projections $W_{c}\to W$. This map will be called $\gamma$.
Lemma 3.4.
The map $\psi:\mathrm{ncS}(V)_{c}\to\mathrm{ncS}(V_{c})$ is a bijective continuous complex affine nc function with continuous inverse $\gamma$.
Proof.
For the readers convenience and because we will need some of the details later, such as certain specific maps, we give two proofs. Since these are closed nc convex sets in a dual operator space we may use the idea in [11, Proposition 2.2.10] to see that it suffices to check this at the $n$th level, for all $n\in\mathbb{N}$. Since $\psi$ is essentially a restriction of the canonical isomorphism $CB(V,\mathbb{R})_{c}\cong CB(V_{c},\mathbb{C})$, it is a continuous complex nc affine function with continuous inverse. To see that this takes $\mathrm{ncS}(V)_{c}$ onto $\mathrm{ncS}(V_{c})$, note that by [3, Lemma 3.1] a map $u:V_{c}\to M_{n}(\mathbb{C})$ is a complex matrix state if and only if its restriction to $V$ is real UCP. However the real UCP maps $h:V\to M_{n}(\mathbb{C})$ are identifiable with the elements $f+ig\in\mathrm{ncS}(V)_{c}$. To see this, notice that the latter are precisely the $f+ig$ such that $c(f,g)(x)=c(f(x),g(x))$ defines a real matrix state on $V$. Indeed these matrix states are precisely the ones which can be identified (via composition with the canonical identification $c_{n}:M_{n}(\mathbb{C})\to M_{2n}(\mathbb{R})$) with a real UCP map $h:V\to M_{n}(\mathbb{C})$ with $h(x)=f(x)+ig(x)$ for $x\in V$. These identifications are another way of describing the map $\psi$ above, and its inverse.
More detailed proof: First, we check that $\psi$ is well defined. Let $f+ig\in(\mathrm{ncS}(V)_{c})_{N}$ so that $c(f+ig)$ will be a ucp map from $V$ to $M_{2N}(\mathbb{R})$. The complexification of $c(f+ig)$ is a ucp map from $V_{c}$ to $M_{2N}(\mathbb{C})$ given by
$\big{(}c(f+ig)_{c}\big{)}(x+iy)=\begin{bmatrix}f(x)&-g(x)\\ g(x)&f(x)\end{bmatrix}+i\begin{bmatrix}f(y)&-g(y)\\ g(y)&f(y)\end{bmatrix}$
Taking the $r$’th amplification shows that for $0\leq[x_{nm}+iy_{nm}]\in M_{r}(V_{c})$ we have
00 0 $\leq\big{(}c(f+ig)_{c}\big{)}^{(r)}([x_{nm}+iy_{nm}])$
$=\begin{bmatrix}\begin{bmatrix}f(x_{nm})&-g(x_{nm})\\ g(x_{nm})&f(x_{nm})\end{bmatrix}+i\begin{bmatrix}f(y_{nm})&-g(y_{nm})\\ g(y_{nm})&f(y_{nm})\end{bmatrix}\end{bmatrix}$
Compressing the last matrix by $\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$ gives that
00 0 $\leq\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\Big{(}\begin{bmatrix}f(x_{nm})&-g(x_{nm})\\ g(x_{nm})&f(x_{nm})\end{bmatrix}+i\begin{bmatrix}f(y_{nm})&-g(y_{nm})\\ g(x_{nm})&f(y_{nm})\end{bmatrix}\Big{)}\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=[f(x_{nm})-g(y_{nm})+if(y_{nm})+ig(x_{nm})]$
$=\psi(f+ig)^{(M)}([x_{nm}+iy_{nm}])$
or $\psi(f+ig)$ is completely positive. It is unital because $c(f+ig)$ is unital and therefore $\psi$ sends elements of $\mathrm{ncS}(V)_{c}$ to matrix states. $\psi$ is continuous because if $f_{\lambda}+ig_{\lambda}\to f+ig$ in $(K_{n})_{c}$ then for all $x,y\in V$ we will have $f_{\lambda}(x)\to f(x)$ and $g_{\lambda}(y)\to y$. So, $\psi(f_{\lambda}+ig_{\lambda})(x+iy)\to\psi(f+ig)$ which means $\lim_{\lambda}\psi(f_{\lambda}+ig_{\lambda})=\psi(f+ig)$.
Note that $\psi$ is graded because it sends certain elements of $M_{n}(CB(V,\mathbb{R})_{c})$ to elements of $M_{n}(CB(V_{c},\mathbb{C}))$. To verify conditions $(2)$ and $(3)$ we show that for any
$f+ig=[f_{nm}+ig_{nm}]\in M_{K}(CB(V,\mathbb{R})_{c})$
and
$\beta=[w_{nm}+iz_{nm}]\in M_{L,K}(\mathbb{C})$
we have that $\psi\big{(}\beta(f+ig)\big{)}=\beta\psi(f+ig)$. To see this, first compute
$[w_{nm}+iz_{nm}][f_{nm}+ig_{nm}]$ $=[\sum_{k}w_{nk}f_{km}-z_{nk}g_{km}+i(z_{nk}f_{km}+w_{nk}g_{km})]$
So, $\psi$ of the above matrix evaluated at $x+iy\in V_{c}$ will be
$[\sum_{n}$ $w_{nk}f_{km}(x)-z_{nk}g_{km}(x)-z_{nk}f_{km}(y)-w_{nk}g_{km}(y)$
$+iw_{nk}f_{km}(y)-iz_{nk}g_{km}(y))+iz_{nk}f_{km}(x)+iw_{nk}g_{km}(x)]$
whereas $[w_{nm}+iz_{nm}](\psi([f_{nm}+ig_{nm}])(x+iy))$ will be
$[w_{nm}+iz_{nm}][f_{nm}(x)-g_{nm}(y)+if_{nm}(y)+ig_{nm}(x)]$
$=[\sum_{n}w_{nk}f_{km}(x)-w_{nk}g_{km}(y)+iw_{nk}f_{km}(y)+iw_{nk}g_{km}(x)$
$+iz_{nk}f_{km}(x)-iz_{nk}g_{km}(y)-z_{nk}f_{km}(y)-z_{nk}g_{km}(x)]$
which are the same.
It follows that $\psi$ will be a complex affine map. The inverse is $\gamma$ since for $\omega\in\mathrm{ncS}(K_{c})$ and $x,y\in V$ we have
$\psi(\text{Re }\omega+i\text{Im }\omega)(x+iy)$ $=\text{Re }\omega(x)-\text{Im }\omega(y)+i\text{Re }\omega(y)+i\text{Im }\omega(x)$
$=(\text{Re }\omega+i\text{Im }\omega)(x+iy)$
$=\omega(x+iy)$
and conversely for $f+ig\in\mathrm{ncS}(V)_{c}$ we get
$\gamma(\psi(f+ig))$ $=\text{Re }\psi(f+ig)+i\text{Im }\psi(f+ig)$
$=f+ig.$
The inverse is a well defined map because Re $\text{Re }\omega+i\text{Im }\omega$ is in $\mathrm{ncS}(V)_{c}$. Indeed $c(\text{Re }\omega+i\text{Im }\omega)$ is ucp because for $M\in\mathbb{N}$ and $0\leq[x_{nm}]\in M_{K}(V)$ we have
$c(\text{Re }\omega+i\text{Im }\omega)^{(K)}([x_{nm}])$ $=\begin{bmatrix}\text{Re }(\omega(x_{nm}))&-\text{Im }(\omega(x_{nm}))\\ \text{Im }(\omega(x_{nm}))&\text{Re }(\omega(x_{nm}))\end{bmatrix}\in M_{2K}(\mathbb{R})$
and the latter is positive if and only if $[\text{Re }(\omega(x_{nm}))+i\text{Im }(\omega(x_{nm}))]=[\omega(x_{nm})]$ is positive. However, this is the $K$’th amplification applied to $[x_{nm}]$ which is positive. Finally, the inverse is continuous because $\psi$ restricted to any $(\mathrm{ncS}(V)_{c})_{N}$ is a bijective continuous map with compact domain. ∎
Example 3.5.
Consider the quaternions $\mathbb{H}$ as a real $C^{*}$-algebra. The state space of $\mathbb{H}$ is trivial, a singleton containing only the map $(a+ib+jc+kd\mapsto a)$. However the non-commutative state space at the higher levels make up for this deficit, and is much more interesting. The complexification of $\mathbb{H}$ is $M_{2}(\mathbb{C})$. So the first layer of $\mathrm{ncS}(\mathbb{H}_{c})\cong\mathrm{ncS}(M_{2}(\mathbb{C}))$ is all states on $M_{2}(\mathbb{C})$, which correspond to positive $2\times 2$ trace one matrices – i.e. the first layer is affine isomorphic to the Bloch sphere. By Lemma 3.4 we have $\mathrm{ncS}(\mathbb{H}_{c})\cong\mathrm{ncS}(\mathbb{H})_{c}$ and so through complexification the first layer of $\mathrm{ncS}(\mathbb{H})$ went from having a single element to containing a three dimensional ball’s worth of elements.
Next, for $K$ real compact nc convex we consider $A(K)_{c}$ as a complexification of an operator system versus $A(K_{c})$ as a complex operator system. Let $f,g:K\to\mathcal{M}(\mathbb{R})$ be real affine maps and $x+iy\in M_{N}(K_{c})$. Define the map $\Psi:A(K)_{c}\to A(K_{c})$ by
$\Psi(f+ig)(x+iy)=u_{n}^{*}\big{(}f(c(x+iy))+ig(c(x+iy))\big{)}u_{n}.$
We will see this has inverse $\Gamma:A(K_{c})\to A(K)_{c}$ taking $\omega\in A(K_{c})$ to Re $\text{Re }\omega+i\text{Im }\omega$.
Lemma 3.6.
For real nc convex sets $K,L$, every real nc affine map $f:K\to L$ has a unique complex nc affine extension $f_{c}:K_{c}\to L_{c}$. If $L$ is complex nc convex there is a unique complex nc affine extension $K_{c}\to L$. These extensions are continuous if $f$ is continuous.
In particular, every real nc affine isomorphism (resp. homeomorphism) $f:K\to L$ has a unique complex nc affine bijective (resp. homeomorphism) extension $f_{c}:K_{c}\to L_{c}$.
Proof.
Define $f_{c}(x+iy)=u^{*}f(c(x,y))u$ if $x+iy\in(K_{c})_{n}$, where $u=u_{n}$ is the isometry above ($u_{n}=1/\sqrt{2}[I_{n}\;-iI_{n}]^{\mathsf{T}}$). Note that $c(\beta)u_{n}=u_{m}\beta$ for $\beta\in M_{m,n}(\mathbb{C})$. Then
$f_{c}(\beta^{*}(x+iy)\beta)=u^{*}f(c(\beta^{*}(x+iy)\beta)u=u^{*}f(c(\beta^{*})c(x,y)c(\beta))u.$
Since $c(\beta^{*})=c(\beta)^{\mathsf{T}}$, and $c(\beta)u=u\beta$, we obtain $f_{c}(\beta^{*}(x+iy)\beta)=\beta^{*}f_{c}(x+iy)\beta$. So $f_{c}$ is affine. A similar argument works if $L$ is complex nc affine. If $f$ is continuous and $x_{n}+iy_{n}\to x+iy$ in $(K_{c})_{n}$ then $x_{n}\to x,y_{n}\to y,c(x_{n},y_{n})\to c(x,y)$. So it is clear from the formula at the start of the proof that $f_{c}(x_{n}+iy_{n})\to f_{c}(x+iy)$, hence $f_{c}$ is continuous.
The uniqueness is clear from the above and the relation $f_{c}(x+iy)=f_{c}(u^{*}c(x,y)u)$. The isomorphism case evidently follows. ∎
Lemma 3.7.
If $f:K\to L$ is a one-to-one continuous nc affine map between closed nc convex sets, and if $K$ is compact, then $f_{n}$ is a homeomorphism onto its (compact) range for all $n$, and $f(K)$ is a compact nc convex set.
Proof.
A continuous one-to-one map on a compact space is a homeomorphism onto its compact range. ∎
We call $f$ in the last result a nc topological affine embedding. Note that if $f$ is a real nc topological affine embedding, then $f_{c}$ is one-to-one and is a complex nc topological affine embedding.
For a real compact nc convex set $K$ there is a canonical map $\epsilon:A(K)\to A_{\mathbb{C}}(K_{c})$ defined by $\epsilon(f)=f_{c}$, where $f_{c}$ is as above.
Lemma 3.8.
The map $\Psi:A(K)_{c}\to A(K_{c})$ is a ucoi with inverse $\Gamma$. Indeed $A(K)$ may be identified with the fixed points of the period 2 conjugate linear complete order automorphism $a_{\theta}$ of $A(K_{c})$ defined by $a_{\theta}(f)=\theta_{\mathbb{C}}\circ f\circ\theta_{K}$, where $\theta_{K}$ is as defined after Theorem 3.1.
Proof.
For the same reason as before, and also to exhibit a complementary result, we give two proofs. Since $\theta_{K}$ is affine and continuous it follows that $a_{\theta}:A(K_{c})\to A(K_{c})$. Since $\theta_{K}$ is period 2, so clearly is $a_{\theta}$ too. Clearly $a_{\theta}$ is unital, and it is not hard to see that it is conjugate linear since $\theta_{\mathbb{C}}$ is conjugate linear: for example if $f\in A(K_{c})$ then
$a_{\theta}(if)(k_{1}+ik_{2})=\theta_{\mathbb{C}}((if)(k_{1}-ik_{2}))=ia_{\theta}(f)(k_{1}+ik_{2}).$
For $x,y\in M_{n}(\mathbb{R})$ we have
$\theta_{\mathbb{C}}((x+iy)^{*})=\theta_{\mathbb{C}}(x^{*}-iy^{*})=x^{*}+iy^{*},$
which equals $(\theta_{\mathbb{C}}(x+iy))^{*}=(x-iy)^{*}$. Thus $a_{\theta}(f^{*})=a_{\theta}(f)^{*}$, since for example
$a_{\theta}(f^{*})(k_{1}+ik_{2})=\theta_{\mathbb{C}}(f^{*}((k_{1}-ik_{2}))=\theta_{\mathbb{C}}(f(k_{1}-ik_{2})^{*}).$
If $f\in M_{n}(A(K))^{+}=A(K,M_{n})^{+}$ then
$[a_{\theta}(f_{ij})(k)]=[\theta_{\mathbb{C}}(f_{ij}(\theta_{K}(k)))],\qquad k\in K_{c}.$
Since $\theta_{\mathbb{C}}$ and its amplifications are completely positive, and $f(K)\geq 0$, we have
$[a_{\theta}(f_{ij})(k)]\geq 0,$
so that $a_{\theta}$ is completely positive.
The fixed points of $a_{\theta}$ clearly include $\epsilon(A(K))$. Indeed $a_{\theta}(f_{c})=\theta_{\mathbb{C}}\circ f_{c}\circ\theta_{K}$ is an affine extension of $f$ and so $a_{\theta}(f_{c})=f_{c}$ by the uniqueness in Lemma 3.6. Conversely, suppose that $a_{\theta}(g)=g$ for $g\in A(K_{c})$. Then $\theta_{\mathbb{C}}(g(\iota(k)))=g(\iota(k))$ for $k\in K$. Thus $g(\iota(k))\in\iota(K)$. Let $f=Re\,g_{|K}$, a real nc affine map on $K$. Then $g=f_{c}=\epsilon(f)$ by the uniqueness in Lemma 3.6, since these are both nc affine extensions of $f$.
More detailed proof: The function $\Psi$ is complex linear. We show that this map is well defined. First, $\Psi(f+ig)$ will be continuous because $f,g$ are continuous, and $\Psi(f+ig)$ is graded because $f,g$ are graded. Now, notice that for matrices $a,b\in M_{N}(\mathbb{R})$ we have
$u_{n}^{*}c(a+ib)^{\mathsf{T}}=(a+ib)^{*}u_{n}^{*}.$
Similarly $c(a+ib)u_{n}=u_{n}(a+ib)$. Let $x+iy\in M_{N}(K_{c})$ and $a,b\in M_{N,k}(\mathbb{R})$ such that $a+ib$ is an isometry, then using facts about the function $c$ in 3.1 and that $f,g$ are real affine gives
$\Psi(f+ig)$ $((a+ib)^{*}(x+iy)(a+ib))$
$=u_{n}^{*}(f(c(a+ib)^{\mathsf{T}}c(x+iy)c(a+ib))$
$\quad+ig(c(a+ib)^{\mathsf{T}}c(x+iy)c(a+ib)))u_{n}$
$=(a+ib)^{*}\big{(}\Psi(f+ig)(x+iy)\big{)}(a+ib).$
Therefore $\Psi(f+ig)$ preserves compressions. A similar proof shows that it preserves direct sums and therefore $\Psi(f+ig)$ is affine. If $f+ig$ is positive in $A(K)_{c}$ then $c(f+ig)$ is positive in $M_{2}(A(K))$, or for any $k\in K$ we have $c(f+ig)(k)\geq 0$. Compressing this matrix by $u_{n}$ gives that $f(k)+ig(k)\geq 0$ for all $k\in K$. From this we see that for any $x+iy\in K_{c}$ we have $\Psi(f+ig)(x+iy)\geq 0$ as we are just taking the adjoint of a positive matrix. Therefore, $\Psi$ is positive and a similar proof shows that our map is completely positive. The unit of $A(K)_{c}$ is $01+i$ 0 where $1$ is the constant function on $K$. $\Psi(1+i0)(x+iy)$ is $1_{N}$ and so $\Psi$ is unital.
To show that $\Psi$ has inverse $\Gamma$ we see that for bounded $\omega:K_{c}\to\mathcal{M}(\mathbb{C})$ and $x+iy\in(K_{c})_{n}$ we have
$\omega(x+iy)$ $=\omega(u_{n}^{*}c(x+iy)u_{n})$
$=u_{n}^{*}\omega(c(x+iy))u_{n}$
$=\Psi((\text{Re }\circ\omega)+i(\text{Im }\circ\omega))(x+iy)=\Psi(\Gamma(\omega))(x+iy).$
Conversely for $f+ig\in A(K)_{c}$ we have
$\Gamma(\Psi(f+ig))$ $=\text{Re }\Psi(f+ig)+i\text{Im }\Psi(f+ig)$
$=f+ig$
where the last equality comes for instance from the fact that
Re $\text{Re }\Psi(f+ig)(x)$ $=\text{Re }\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\big{(}f(x\oplus x)+ig(x\oplus x)\big{)}\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=\text{Re }(f(x)+ig(x))=f(x).$
Finally, we need to show that $\Gamma$ is ucp. It is unital because $\Psi$ is unital. If $[\omega_{nm}]\in M_{N}(A(K_{c}))$ is positive then it maps all $x+iy\in K_{c}$ to positive matrices. So, for $x\in K$ we have
$c([\text{Re }\circ\omega_{nm}+i\text{Im }\circ\omega_{nm}])(x)=c([\text{Re }(\omega_{nm}(x))+i\text{Im }(\omega_{nm}(x))])=c([\omega_{nm}(x)])$
Now $[\omega_{nm}(x)]$ is positive in $M_{N}(\mathbb{C})$ and so $c([\omega_{nm}(x)]$ is positive in $M_{2N}(\mathbb{R})$. Therefore, $c([\Gamma(\omega_{nm})])$ is a positive element of $M_{2}(A(K))$ which by definition means $[\Gamma(\omega_{nm})]$ is positive in $A(K)_{c}$.
4. Real Categorical Duality
For any real compact nc convex set $K$, $A(K)$ will be a real operator system by 2.4. On the other hand, given a real operator system $V$, $\mathrm{ncS}(V)$ will be a real compact nc convex set. We have the following duality extending the complex case in Theorem 3.2.2 of [11]:
Theorem 4.1.
Let K be a (real or complex) compact nc convex set, then $K\cong\mathrm{ncS}(A(K))$ via the complex affine homeomorphism $\Lambda:K\to\mathrm{ncS}(A(K))$ where
$\Lambda(x)(\varphi)=\varphi(x)$
for $x\in K$ and $\varphi\in\mathrm{ncS}(A(K))$.
Conversely we have (extending Theorem 3.2.3 in [11]):
Theorem 4.2.
Let $V$ be a closed (real or complex) operator system. For $v\in V$ define the function $\hat{v}:\mathrm{ncS}(V)\to\mathcal{M}(\mathbb{F})$ by
$\hat{v}(\varphi)=\varphi(v)$
for $\varphi\in\mathrm{ncS}(V)$. This map is a continuous nc affine function. The map ^$\hat{\phantom{v}}:V\to A(\mathrm{ncS}(V))$ taking $v$ to $\hat{v}$ is a ucoi.
We first prove Theorem 4.1 in the real case. For a real compact nc convex set, we have that $K_{c}$ is a complex nc convex set and therefore isomorphic to $\mathrm{ncS}(A(K_{c}))$ which in turn is isomorphic to $\mathrm{ncS}(A(K))_{c}$. We have the embedding $\iota$ of our real nc convex sets into their complexification, so we just need to make sure $\iota(K)$ is mapped onto $\iota(\mathrm{ncS}(A(K))$ through the above isomorphisms. Or, the following diagram commutes
[…]
Here, $\Psi^{*}:\mathrm{ncS}(A(K_{c}))$ is defined by $\Psi^{*}(f)(x+iy)=f(\Psi(x+iy))$. So now we just diagram chase. Let $k\in K_{N}$, then going to the right, for $f+ig\in A(K)_{c}$ we have
$\psi(\Lambda(k)+i0)(f+ig)$ $=\Lambda(k)(f)+i\Lambda(k)(g)=f(k)+ig(k).$
Going up we have:
$\Psi^{*}(\Lambda(k+i0))(f+ig)$ $=\Lambda(k+i0)(\Psi(f+ig))$
$=\Psi(f+ig)(k+i0)$
$=\frac{1}{2}\begin{bmatrix}1_{N}&i\cdot 1_{N}\end{bmatrix}\big{(}f(c(k+i0))+ig(c(k+i0))\big{)}\begin{bmatrix}1_{N}\\ -i\cdot 1_{N}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}1_{N}&i\cdot 1_{N}\end{bmatrix}\big{(}f(k\oplus k)+ig(k\oplus k)\big{)}\begin{bmatrix}1_{N}\\ -i\cdot 1_{N}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}1_{N}&i\cdot 1_{N}\end{bmatrix}\big{(}f(k)\oplus f(k)+ig(k)\oplus g(k)\big{)}\begin{bmatrix}1_{N}\\ -i\cdot 1_{N}\end{bmatrix}$
$=f(k)+ig(k).$
Therefore, starting at the bottom left and going clockwise, is the same as going right and then anticlockwise in the diagram. That is, the diagram commutes, resulting in a nc affine homeomorphism between $K$ and $\mathrm{ncS}(A(K))$.
Remark. Alternatively, the the proof of Theorem 3.2.2 of [11] works in the real case. For the readers convenience we give an alternative proof of surjectivity of the map there. Namely, if $\varphi\in\mathrm{ncS}(A(K))$ then $\iota(\varphi)\in\mathrm{ncS}(A(K))_{c}=\mathrm{ncS}_{\mathbb{C}}(A_{\mathbb{C}}(K_{c}))$. (Indeed $\varphi_{c}$ is a matrix state of $A(K)_{c}$, and thus gives a matrix state of $A_{\mathbb{C}}(K_{c})$ by composition with the canonical map $A_{\mathbb{C}}(K_{c})\to A(K)_{c}$.) Thus by Theorem 3.2.2 of [11] there exists $x+iy\in K_{c}$ with $\theta(x+iy)=\iota(\varphi)$. That is, for all $f\in A_{\mathbb{C}}(K_{c})$ we have
$f(x+iy)=\iota(\varphi)(f)=\varphi({\rm Re}f_{|K})+i\varphi({\rm Im}f_{|K}).$
Here ${\rm Re},{\rm Im}$ here are the real affine functions coming from Lemma 2.5. In particular, replacing $f$ by $\epsilon(f)=f_{c}$ for $f\in A(K)$, taking real parts, and remembering that $x\in K_{n}$, we have $f(x)={\rm Re}(f(x)+if(y))=\varphi(f)$ for $f\in A(K).$ That is $\Lambda(x)=\varphi$, so that $\Lambda$ is surjective.
The proof of 4.2 in the real case is similar. We want to show the following diagram commutes
[…]
To show this, let $v\in V$. Going to the right we have
$\Psi(\hat{v}+i0)(x+iy)$ $=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\big{(}\hat{v}(c(x+iy)\big{)}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\big{(}c(x(v)+iy(v))\big{)}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=x(v)+iy(v)$
and going up we get
$\psi^{*}(\widehat{v+i0})(x+iy)$ $=(\widehat{v+i0})(\psi(x+iy))$
$=\psi(x+iy)(v+i0)$
$=x(v)+iy(v)$
It follows as in [11] that the categories NCConvR (real compact nc convex sets and continuous affine nc maps) and OpSyR (real operator systems and ucp maps) are dually equivalent via a contravariant functor. Thus for example $f\mapsto f\circ\tau:A(K_{2})\to A(K_{1})$, composition of $f\in A(K_{2})$ with a continuous affine nc map $\tau:K_{1}\to K_{2}$, is ucp. For compact nc convex sets $K$ and $L$ we have $A(K)$ and $A(L)$ unitally complete order isomorphic if and only if $K$ and $L$ are affinely homeomorphic. Hence two operator systems are isomorphic if and only if their nc state spaces are affinely homeomorphic.
We now characterize the real compact nc convex sets which correspond to real operator systems which are the selfadjoint part of another operator system.
Corollary 4.3.
A real compact nc set $K$ corresponds under the duality above to a real operator system with trivial involution if and only if every $k\in K$ is symmetric (that is, $k=k^{\mathsf{T}}$).
Proof.
For ucp $\varphi:V\to M_{n}$ we have
$\varphi(x)^{\mathsf{T}}=\varphi(x^{*})\qquad x\in V.$
Thus if $V$ has trivial involution then $K=\mathrm{ncS}(V)$ is symmetric. Similarly, if the latter is symmetric then $\varphi(x^{*})=\varphi(x)$ for all such $\varphi$ and $x\in V$, so that $x=x^{*}$. ∎
We shall call real compact nc sets satisfying the condition in the last result symmetric.
5. Bipolar Theorem
5.1. Real Bipolar Theorem
Let $E$ be a real dual operator space and $K$ a real nc set over $E$. The polar of $K$ is a real nc convex set over $E^{*}$ defined by
$K_{n}^{\circ}=\{\varphi\in M_{n}(E^{*}):{\mathfrak{R}}\,\varphi^{(m)}(v)\leq 1_{nm}\text{ for all }v\in K_{m},m\leq\kappa\}$
The set $K^{\circ}=\bigsqcup_{n}K_{n}^{\circ}$ is a closed real nc convex set. This definition is the same as in the complex case, except that Effros and Winkler only consider finite $n,m$ in their definition. Note though that if ${\mathfrak{R}}\,\varphi^{(m)}(v)\leq 1_{nm}$ for all finite $m$ and $v\in K_{m},$ then it is easy to argue that we have the same relation for infinite $m$. Thus we may take $m<\infty$ in the definition above. Hence the definition makes sense and produces a closed real nc convex set even if $K$ is only a matrix convex set.
Proposition 5.1.
Let $K$ be a real nc set (or real matrix set) in the real dual operator space $E$. Then $(K^{\circ})_{c}\cong(K_{c})^{\circ}$ via the same maps as in Lemma 3.4.
Proof.
Of course $K_{c}$ is a complex nc convex set containing $0_{E_{c}}\in E_{c}$. Let $\gamma:(K_{c})^{\circ}\to(K^{\circ})_{c}$ taking $\omega$ to $\text{Re }\omega_{|E}+i\text{Im }\omega_{|E}$, which will have inverse $\psi$. To see that $\gamma$ maps into $(K^{\circ})_{c}$, let $\omega\in(K_{c})^{\circ}_{n}$ and $v\in K_{m}$. We have
${\mathfrak{R}}\,c(\gamma(\omega))^{(m)}(v)$ $={\mathfrak{R}}\,c(\text{Re }\omega^{(m)}(v+i0)+i\text{Im }\omega^{(m)}(v+i0))$
$={\mathfrak{R}}\,c(\omega^{(m)}(v+i0))$
$=c({\mathfrak{R}}\,\omega^{(m)}(v+i0)).$
We also have
$c({\mathfrak{R}}\,\omega^{(m)}(v+i0))\leq 1_{m\cdot 2n}\iff{\mathfrak{R}}\,\omega^{(m)}(v+i0)\leq 1_{mn}.$
However if $v\in K$ then $v+i0\in K_{c}$, and so this is true. So $\gamma$ maps into $(K^{\circ})_{c}$. We saw in Lemma 3.4 that $\gamma$ is a bijective continuous affine map with continuous inverse $\psi$. Hence we will be done if $\psi((K^{\circ})_{c})\subseteq(K^{\circ})_{c}$.
The inverse map $\psi:(K^{\circ})_{c}\to(K_{c})^{\circ}$ takes $f+ig\in((K^{\circ})_{c})_{N}$ and $x+iy\in E_{c}$ to
$\psi(f+ig)(x+iy)=f(x)-g(y)+if(y)+ig(x).$
To show this indeed maps into $(K_{c})^{\circ}$, suppose that $x+iy=[x_{nm}+iy_{nm}]\in(K_{c})_{M}$, so that $c(x+iy)\in K_{2M}$. We want to show that
${\mathfrak{R}}\,[f(x_{nm})-g(y_{nm})+if(y_{nm})+ig(x_{nm})]\leq 1_{NM}.$
However, as shown in 3.4 we have
${\mathfrak{R}}\,[$ $f(x_{nm})-g(y_{nm})+if(y_{nm})+ig(x_{nm})]$
$={\mathfrak{R}}\,u_{N}^{*}\Big{(}\begin{bmatrix}f(x_{nm})&-g(x_{nm})\\ g(x_{nm})&f(x_{nm})\end{bmatrix}+i\begin{bmatrix}f(y_{nm})&-g(y_{nm})\\ g(y_{nm})&f(y_{nm})\end{bmatrix}\Big{)}u_{N}$
$={\mathfrak{R}}\,u_{N}^{*}\big{(}c(f+ig)^{(M)}(x)+ic(f+ig)^{(M)}(y)\big{)}u_{N}$
$=u_{N}^{*}{\mathfrak{R}}\,\big{(}c(f+ig)^{(M)}(x)+ic(f+ig)^{(M)}(y)\big{)}u_{N}.$
Because $u_{n}$ is an isometry we see that the latter is $\leq 1_{NM}$ as for $f+ig\in(K^{\circ})_{c}$ and $x+iy\in K_{c}$ we have that
${\mathfrak{R}}\,c(f+ig)^{(2N)}(c(x+iy))={\mathfrak{R}}\,\begin{bmatrix}c(f+ig)^{(N)}(x)&-c(f+ig)^{(N)}(y)\\ c(f+ig)^{(N)}(y)&c(f+ig)^{(N)}(x)\end{bmatrix}\leq 1_{4NM}.$
This completes the proof. ∎
Theorem 5.2 (Bipolar Theorem).
Let $K\subseteq E$ be a closed real or convex nc convex set containing $0_{E}\in E$. Then $K^{\circ\circ}\cong K$.
Proof.
Clearly $K\subseteq K^{\circ\circ}$. If $x\in(K^{\circ\circ})_{n}\setminus K_{n}$ then by Theorem 3.3 there exists a normal completely bounded map $\varphi:E\to M_{n}(\mathbb{F})$ such that ${\mathfrak{R}}\,\varphi_{n}(x)\not\leq 1_{n}\otimes 1_{n}$ but for all $p$ and $k\in K_{p}$ we have ${\mathfrak{R}}\,\varphi_{p}(k)\leq 1_{p}\otimes 1_{n}$. Then $\varphi\in(K^{\circ})_{n}$, and we obtain the contradiction $\varphi_{n}(x)\leq 1_{n}\otimes 1_{n}$. So $K=K^{\circ\circ}$. ∎
Effros and Winkler’s application of the bipolar theorem in [13, Section 5] essentially works for us too. That is a weakly compact nc convex set $L$ with $L_{1}=K$ is sandwiched between minimal and maximal nc convex sets which are $K$ at level 1. Up to translation, the maximal one is defined as the prepolar of the ‘minimal one’ of $L^{\circ}$. In the real case, though, this is a bit more helpful under the restriction that the compact nc convex set corresponds to an operator system with trivial involution. The reader can see this by analyzing the application in the case of the noncommutative state space of the quaternions. We will discuss this a little further at the end of the next section.
6. Max and Min nc convex sets
A classical compact convex set $K$ may be turned into a compact nc convex set using the fact that an operator system can be given a minimum and maximum operator system structure OMIN and OMAX [23, 31], and then employing the categorical duality between compact nc convex sets and operator systems. Thus we define ${\rm Min}(K)$ and ${\rm Max}(K)$ by $A({\rm Min}(K))={\rm OMIN}(A(K))$ and $A({\rm Max}(K))={\rm OMAX}(A(K))$. For a complex operator system $V$ we thus get the ‘minimal and maximal nc convex sets’ as exactly the nc convex sets corresponding to OMIN$(V)$ and OMAX$(V)$. Because of the issues in the real case with OMIN and OMAX (which restricts us to the real operator systems with trivial involution), and also in order to establish some complementary results, we will proceed slightly differently.
By a real function system (in Kadison’s sense, so that there is ‘no involution’, that is there is a trivial involution) we mean (concretely) a unital subspace of $C_{\mathbb{R}}(K)$ for a compact Hausdorff $K$, or abstractly (via Kadison’s theorem), a (real) ordered vector space with an archimedean order unit.
Similarly a complex function system is (concretely) a unital selfadjoint subspace of $C_{\mathbb{C}}(K)$ for compact $K$, or abstractly [22, 29], a complex *-vector space which is ordered (i.e. with a proper selfadjoint cone $E^{+}$), and possesses an archimedean order unit.
These form categories, with the morphisms being unital selfadjoint positive maps, or equivalently (by basic results in e.g. [21, Section 2]) unital selfadjoint contractions.
Proposition 6.1.
The category of complex function systems is equivalent to the category of real function systems. Moreover every real function system has a unique reasonable function system complexification, and every complex function system has unique real structure, that is, is the complexification of an essentially unique real function system.
Proof.
The ‘function system complexification’ of a real function system $S$ is unique and reasonable (we assume that the embedding of $S$ into the function system complexification is as real valued functions). This complexification is the so-called Taylor complexification (see e.g. [19]).
Conversely, every complex function system $V$ is a reasonable complexification of a real function system. Indeed suppose that $V$ is a unital selfadjoint subspace of $C(K,\mathbb{C})$. Then $V_{\rm sa}$ is a real function system with the inherited cone (for example it is clearly an ordered real space with archimedean order unit $1_{V}$). Moreover $V$ is a reasonable complexification of $V_{\rm sa}$.
Clearly the above defines two functors between the categories. Notice that a unital positive map $T:V\to W$ in the complex category is completely positive, and selfadjoint, so $T(V_{\rm sa})\subseteq W_{\rm sa}$. Thus it is clear that the category of complex function systems is isomorphic to the category of real function systems.
Any system $S$ whose complexification is $V$ is the fixed points for some period 2 conjugate linear unital order isomorphism $u:V\to V$. Note that $u$ is selfadjoint since it is positive. Let $w=u_{|V_{\rm sa}}$, which is a period 2 order automorphism of $V_{\rm sa}$. Then $u\circ w_{c}$ is a period 2 conjugate linear unital order isomorphism $V\to V$ whose fixed points are exactly $V_{\rm sa}$. That is, $u(w_{c}(v))=\bar{v}$ for $v\in V$. Thus up to the unital order isomorphism $w_{c}$, the complexification of $S$ can be identified with the complexification of $V_{\rm sa}$. ∎
Remark. Of course the analogue for operator systems of most of the assertions in the last result is (badly) false. In particular a complex operator system $V$ need not be a reasonable complexification of $V_{\rm sa}$. There is however Lemma 2.6. There is also a one-to-one real continuous nc affine map ${\rm ncS}_{\mathbb{C}}(V)\to{\rm ncS}_{\mathbb{R}}(V_{\rm sa})$, taking $\varphi\mapsto\text{Re }\,\varphi$. Since Re is completely contractive this map is well defined. It is one-to-one since Re $\varphi=0$ implies that Re $(i\varphi)=0$. It is surjective if $V={\rm OMAX}(V)$ for example.
We now consider the nc convex sets canonically associated with the function system and its complexification.
Let $K$ be a classical compact convex set in a real dual Banach space $E$. Since the beginnings of the subject of matrix and nc convexity, authors have shown that in the complex case there is a smallest and largest matrix/nc convex set which agrees at the first level with $K$ (see e.g. [13, Section 5], or [20, Section 1.2.3] and references therein). It seems to us that these sets depend on the particular embedding of $K$ into a LCTVS operator space. However in this paper we will define, in both the real and complex case, Min$(K)$ to be the closed nc convex hull in Max$(E)$ of $(K,\emptyset,\emptyset,\cdots)$. This is the smallest compact nc convex set containing $(K,\emptyset,\emptyset,\cdots)$. We remark that our notation conflicts with that in [17, Section 5], who call this max$(K)$ because they want to regard it as the largest compact nc convex in their ordering.
Lemma 6.2.
Let $K$ be a classical compact convex set as above. At the first level ${\rm Min}(K)$ is simply $K$, at the $n$-th level it (that is, $({\rm Min}(K)_{n}$) is the weak* closure in $M_{n}(E)$ of the ordinary convex hull ${\rm co}(C)$ of the set $C$ of terms $a\otimes k$ for a (trace 1 positive selfadjoint) density matrix $a\in M_{n}(\mathbb{F})^{+}$ and $x\in K$.
Proof.
To see this first note that
${\rm co}(C)=\{\sum_{j=1}^{m}\,v_{j}^{*}k_{j}v_{j}:m\in\mathbb{N},k_{j}\in K,v_{j}\in M_{1,n},\sum_{j}\,v_{j}^{*}v_{j}=I_{n}\}.$
Indeed if $x$ is in this weak* closure $W_{n}$ in $M_{n}(E)$, a limit of $x_{t}$, where $x_{t}\in{\rm co}(C)$, and $\beta\in M_{m,n}$ is an isometry, then $\beta^{*}x_{t}\beta\to\beta^{*}x\beta$ weak*. Thus these weak* closures $(W_{n})$ satisfies (3) in the definition of a nc convex set, and similarly it satisfies (2) there, if we also use the fact that $\sum_{i}\,\alpha_{i}x_{i}\alpha_{i}^{*}$ converges weak*. Hence $(W_{n})$ is a nc convex set. It is closed, and hence compact. Indeed we may assume that $K\subseteq{\rm Ball}(E)$, and hence in $M_{n}(E)$ the nc convex combinations are in ${\rm Ball}(M_{n}({\rm Max}(E)))$. Since the latter ball is weak* compact, so is $W$. Clearly then this is the smallest closed nc convex set which agrees at the first level with $K$.
The statement about the nonentangled matrix states follows from the description above and the next lemma (see the line before the lemma). ∎
We remark that the weak* closure is unnecessary if $A(K)$ is finite dimensional and $n<\infty$. For the convex hull of a compact set in a finite dimensional LCTVS is compact, and the set $C$ above is easily seen to be compact in $M_{n}(E)$.
If $K$ is the classical state space of an operator system and $n$ is a cardinal then the weak* closed convex hull $W_{n}$ of the set $C$ defined in Lemma 6.2, may be called the separable (i.e. nonentangled) matrix states of $V$. The lemma asserts that these nonentangled states (for all levels $n$, i.e. $W=(W_{n})$) is the nc convex set Min$(K)$.
We shall see next that this is also exactly the nc/matrix state space of OMIN$(A(K))$.
Lemma 6.3.
Let $\varphi:V\to B(H)$ be a completely positive selfadjoint map on a real operator system. Let $a=\varphi(1)^{\frac{1}{2}}$. Then there exists ucp $\Psi:V\to B(H)$ such that $\varphi=b\Psi(\cdot)b$.
Proof.
Note that $\varphi:V_{c}\to B(H)_{c}$ is completely positive, so by e.g. [10, Lemma 2.2] there exists ucp $\Psi:V_{c}\to B(H)_{c}$ with $\varphi_{c}=b\Psi(\cdot)b$. Inspecting the proof of the last cited result we see that $\Psi(V)\subseteq B(H)$ and $\varphi=b\Psi(\cdot)b$. ∎
Proposition 6.4.
Let $K$ be a classical compact convex set. Then $A({\rm Min}(K))={\rm OMIN}(A(K))$. That is, ${\rm Min}(K)$ is the nc convex set corresponding to ${\rm OMIN}(A(K))$ via the functorial correspondence between compact nc convex sets and operator systems. In particular, ${\rm Min}(K)$ consists of the ucp maps ${\rm OMIN}(A(K))\to M_{n}$, for all $n\leq\kappa$.
Proof.
To prove this note that if $f\in A({\rm Min}(K))$ then clearly $f\in A(K)$. Conversely, suppose that $f\in A(K)$, and $x\in{\rm Min}(K)_{n}$. We may assume that $K$ is a subset of a real dual Banach space such that $f$ extends to a linear continuous $\varphi\in E^{*}$. Indeed we can take $E={\rm Max}(A(K)^{*})$. The restriction of $\varphi_{n}$ to $K_{n}$ defines the desired function from Min$(K)_{n}$ to $M_{n}$ extending $f$. Call this function $f_{n}$, then $f_{n}(R\otimes k)=Rf(k)$ for density matrix $R\in M_{n}^{+}$ and $k\in K$. This defines (the unique) nc affine function $\hat{f}$ in $A({\rm Min}(K))$, which at first level is $f:K\to\mathbb{F}$.
Note that $\|\hat{f}_{n}\|=\|f\|$. To see this note that for $x=\sum_{j}\,v_{j}^{*}k_{j}v_{j}$ with $k_{j}\in K,v_{j}\in M_{1,n}$ and $\sum_{j}\,v_{j}^{*}v_{j}=I_{n}$ we have
$\|\hat{f}_{n}(x)\|=\|\sum_{j}\,v_{j}^{*}f(k_{j})v_{j}\|\leq\max_{j}\,|f(k_{j})|\leq\|f\|.$
By density and continuity we see that $\|\hat{f}_{n}(x)\|\leq\|f\|$ for all $x\in({\rm Min}(K))_{n}$.
This proves that the canonical map $A({\rm Min}(K))\to{\rm OMIN}(A(K))$ is a unital isometry, and hence by a basic property of OMIN it is a completely positive complete contraction. Conversely, suppose that $f=[f_{ij}]\in M_{n}({\rm OMIN}(A(K)))^{+}$, so that $f(k)\geq 0$ for all $k\in K$. Claim: $[\widehat{f_{ij}}]\in M_{n}(A({\rm Min}(K)))^{+}$. Indeed for $x$ as in the last paragraph, we have for an appropriate scalar matrix $V$ that
$[\widehat{f_{ij}}(x)]=[\sum_{j}\,v_{j}^{*}f_{ij}(k_{j})v_{j}]=V^{*}{\rm diag}(f(k_{1}),\cdots,f(k_{n}))V\geq 0.$
Thus $[\widehat{f_{ij}}]\geq 0$, by density and continuity. This proves the Claim, so that the canonical map $A({\rm Min}(K))\to{\rm OMIN}(A(K))$ is a unital complete order isomorphism, hence also a complete isometry. ∎
Most of the last result in the complex case and for $n<\infty$ can also be deduced from an assertion in [23, Theorem 4.8 and Remark 4.5], and is equivalent to that assertion. Indeed we use the above to give a generalization of this result:
Corollary 6.5.
Let $V$ be a real or complex function system, $H$ a real or complex Hilbert space, and $n<\infty$. Any element of $M_{n}({\rm OMIN}(V)^{d})^{+}$, or any completely positive selfadjoint map ${\rm OMIN}(V)\to B(H)$, is a point weak* limit of a uniformly bounded net of maps of the form $\sum_{j}\,v_{j}^{*}\varphi_{j}v_{j}=\sum_{j}\,(v_{j}^{*}v_{j})\otimes\varphi_{j}$, for (scalar valued) states $\varphi_{j}$ on $V$ and row vectors $v_{i}$ with real or complex entries and $\sum_{j}\,v_{j}^{*}v_{j}$ strongly convergent.
Proof.
Let $\varphi:V\to B(H)$ be completely positive and selfadjoint. Let $b=\varphi(1)^{\frac{1}{2}}$. By Lemma 6.3 there is a UCP $\Psi:V\to B(H)$ such that $\varphi=b\Psi(\cdot)b$. By Lemma 6.2 and its proof we have that $\Psi$ is a point weak* limit of maps of the form $\sum_{i}\,v_{i}^{*}\varphi_{j}v_{j}$ for states $\varphi_{j}$ and with $\sum_{i}\,v_{i}^{*}v_{j}=1$. Hence $\varphi$ is a point weak* limit of maps of the form $\sum_{j}\,bv_{j}^{*}\,\varphi_{j}(\cdot)\,v_{j}b$. Note that $\|\sum_{j}\,bv_{j}^{*}v_{j}b\|=\|\varphi(1)\|$, so the net is uniformly bounded by $\|\varphi(1)\|$. ∎
Lemma 6.6.
Let $K$ be a classical compact convex set (in a real dual Banach space). Then ${\rm Min}_{\mathbb{R}}(K)_{c}={\rm Min}_{\mathbb{C}}(K)$ as nc convex sets, with both equaling $K$ at the first level.
Proof.
Complexifying the relation $A_{\mathbb{R}}({\rm Min}_{\mathbb{R}}(K))={\rm OMIN}_{\mathbb{R}}(A(K)),$ we have
$A_{\mathbb{C}}(({\rm Min}_{\mathbb{R}}(K))_{c})={\rm OMIN}_{\mathbb{R}}(A(K))_{c}={\rm OMIN}_{\mathbb{C}}(A_{\mathbb{C}}(K)).$
In the last equality we used [6, Proposition 9.18] taking $V={\rm OMIN}(A_{\mathbb{R}}(K))$, so that $V_{c}=A_{\mathbb{C}}(K)$ (note that $A_{\mathbb{R}}(K)$ has the trivial involution at the first level). However ${\rm OMIN}_{\mathbb{C}}(A_{\mathbb{C}}(K))=A_{\mathbb{C}}({\rm Min}_{\mathbb{C}}(K))$ by the complex case of Proposition 6.4. It follows by the functorial correspondence between compact nc convex sets and operator systems that $({\rm Min}_{\mathbb{R}}(K))_{c}$ is the complex closed nc convex hull $\overline{{\rm co}}_{\mathbb{C}}(K,\emptyset,\cdots)={\rm Min}_{\mathbb{C}}(K,\emptyset,\cdots)$. That is, ${\rm Min}_{\mathbb{R}}(K)_{c}={\rm Min}_{\mathbb{C}}(K)$. ∎
An OMIN (resp. OMAX) operator system is just a (real or complex) function system with OMIN (resp. OMAX) operator system structure, or equivalently equals OMIN$(A(K))$ (resp. OMAX$(A(K))$) for a classical closed convex set $K$. As discovered in [6, Section 9], an AOU space with nontrivial involution cannot be an OMIN (resp. OMAX) operator system.
The maximal quantization ${\rm Max}(K)$ for us is different in general from the one usually defined by duality or by the bipolar theorem as the prepolar of $({\rm Min}(K))^{\circ}$ (see e.g. [13, End of Section 5]). We define ${\rm Max}(K)$ by the functorial correspondence between compact nc convex sets and operator systems, via $A({\rm Max}(K))={\rm OMAX}(A(K))$. We have that Max$(K)$ is the nc set $(K^{n})$ with $K^{n}$ in $M_{n}(E^{*})$ the set of unital positive (selfadjoint) linear maps $\varphi:A_{\mathbb{F}}(K)\to M_{n}$. This is nc convex and nc compact, indeed by a basic property of OMAX, $(K^{n})$ is the nc matrix state space of ${\rm OMAX}(A(K))$. That is,
Lemma 6.7.
Let $K$ be a classical compact convex set (in a dual Banach space). Then ${\rm Max}(K)={\rm ncS(OMAX}(A(K)))$.
Proposition 6.8.
Let $K$ be a classical compact convex set (in a real dual Banach space). Then ${\rm Max}_{\mathbb{R}}(K)_{c}={\rm Max}_{\mathbb{C}}(K)$ as nc convex sets, with both equaling $K$ at the first level.
Proof.
This is almost identical to the OMIN case. Complexifying the relation $A_{\mathbb{R}}({\rm Max}_{\mathbb{R}}(K))={\rm OMAX}_{\mathbb{R}}(A(K)),$ we have
$A_{\mathbb{C}}(({\rm Max}_{\mathbb{R}}(K))_{c})={\rm OMAX}_{\mathbb{R}}(A(K))_{c}={\rm OMAX}_{\mathbb{C}}(A_{\mathbb{C}}(K)).$
In the last equality we used [6, Proposition 9.18] with $V={\rm OMIN}(A_{\mathbb{R}}(K))$, so that $V_{c}=A_{\mathbb{C}}(K)$. However ${\rm OMAX}_{\mathbb{C}}(A_{\mathbb{C}}(K))=A_{\mathbb{C}}({\rm Max}_{\mathbb{C}}(K))$ by the discussion above the lemma. It follows by the functorial correspondence between compact nc convex sets and operator systems that $({\rm Max}_{\mathbb{R}}(K))_{c}={\rm Max}_{\mathbb{C}}(K)$. ∎
Remark. We saw in the proof above that $({\rm OMAX}_{\mathbb{R}}(A(K)))_{c}={\rm OMAX}_{\mathbb{C}}(A_{\mathbb{C}}(K))$. This is related to the fact that the ‘function system complexification’ of a real function system is unique.
One might think that it is obvious that ${\rm Min}(K)\subseteq{\rm Max}(K)$ since ${\rm Max}(K)$ is a nc convex set containing $(K,\emptyset,\emptyset,\cdots)$. However recall that ${\rm Min}(K)$ naturally ‘lives in’ ${\rm Max}(A(K)^{*})$, being the noncommutative states on OMIN$(A(K))$, while ${\rm Max}(K)$ corresponds to the noncommutative states on OMAX$(A(K))$, and naturally ‘lives in’ Min$(A(K)^{*})$. Of course ${\rm Max}(K)$ are the unital selfadjoint positive maps from OMIN$(A(K))$ into $M_{n}$, while ${\rm Min}(K)$ is the subset of ucp maps into $M_{n}$. This is not a problem for finite $n$, since one may just use the ‘product topology’ (see e.g. 1.6.4 in [5]). For infinite $n$ this suggests that perhaps the ‘ambient LCTVS operator space’ for both is Min$(A(K)^{*})$. In any case, the identity map, which is the adjoint of the canonical ucp map $\Phi:{\rm OMAX}(A(K))\to{\rm OMIN}(A(K))$, yields a canonical nc affine continuous map ${\rm Min}(K)\to{\rm Max}(K)$. It is surjective at the first level of course.
Proposition 6.9.
Let $K$ be a classical compact convex set (in a dual Banach space). The canonical nc affine embedding $\epsilon:{\rm Min}(K)\to{\rm Max}(K)$ is a nc topological affine embedding. That is, $\epsilon$ is a homeomorphism onto its (compact) range for all $n$. Similarly for a closed nc convex set $L$ in a dual operator space $E$ with $L_{1}=K$, with $L$ symmetric if $\mathbb{F}=\mathbb{R}$, there are canonical nc topological affine embeddings Min$(K)\subseteq L\subseteq{\rm Max}(K)$.
Proof.
To see that $\epsilon$ is one-to-one note that if $\varphi:{\rm OMIN}(A(K))\to M_{n}$ satisfies $\varphi\circ\Phi=0$ for $\Phi$ as above, then $\varphi=0$. The statement then follows from Lemma 3.7.
For the last assertion we suppose that $K$ corresponds to an operator system $V$. If $K$ is symmetric $V$ has trivial involution. ‘Level 1’ of $V$ is a function system with state space $K$. The canonical ucp maps
${\rm OMAX}(A(K))\to V=A(L)\to{\rm OMIN}(A(K))$
dualize to give continuous nc affine maps
${\rm Min}(K)\to L\to{\rm Max}(K)\subseteq{\rm Min}(A(K)^{*}.$
As in the last paragraph these maps are one-to-one, and nc topological affine embeddings by Lemma 3.7. ∎
The following is the real version of [11, Theorem 2.5.8]:
Theorem 6.10.
If $K$ is a symmetric real compact nc convex set then the canonical restriction map $\rho:A(K)\to A(K_{1})$ is a contractive unital order isomorphism whose inverse map has norm $\leq 2$.
Proof.
Clearly $\rho$ is contractive, unital and positive. Indeed $\rho$ is one of the nc topological affine embeddings in the last assertion in Proposition 6.9, so is an order isomorphism. Since for $a\in A(K)$ and $x\in K$ we have $a(x)^{*}=a(x)$ the proof of [11, Theorem 2.5.8] involving the polarization identity works to see that $\|a\|\leq 2\|a_{|{K_{1}}}\|$. ∎
Remarks. 1) The first assertions of [11, Theorem 2.5.8] are true for all real symmetric compact nc convex sets by a similar proof.
2) The necessity of the ‘symmetric’ condition here is clear by examining the case of the quaternions.
[…]
Remarks. 1) One cannot however expect analogues of Propositions 6.4 and 6.8 to hold in general for Mink and Maxk. E.g. if $K$ be a compact nc convex set (in a real dual operator space) then often ${\rm Max}_{\mathbb{R},k}(K)_{c}\neq{\rm Max}_{\mathbb{C},k}(K)$. Indeed this fails in general, as do the matching operator system equalities (matching via the functorial correspondence between compact nc convex sets and operator systems). For $k=1$ it fails because of the existence of real entangled states that are complex separable (i.e. nonentangled) (see [9] and [6, Section 9]). For $k>1$ it fails because of the problems with complexifying entanglement breaking maps as seen in [9] and [6, Section 9].
2)  Effros and Winkler’s application of the bipolar theorem discussed at the end of Section 5 is not in full generality necessarily well related to Min and Max as we define them above. Indeed, if $L$ is a closed nc convex set in $E$ with $L_{1}=K$, then one cannot even necessarily expect Min$(K)\subseteq L\subseteq{\rm Max}(K)$. Here by $\subset$’ we mean a canonical nc topological affine embedding. Note that the second inclusion here is certainly not true if $L$ is the nc state space of the quaternions. That Min$(K)\subseteq L$ is usually clear in applications, but may not be very helpful. We can however enlarge the definition of Max, defining it for example for any operator system $V$ to be the closed nc convex set $C$ that at level $n$ corresponds to the unital positive selfadjoint maps $V\to M_{n}$ of norm $\leq 2$ (we need some condition such as the latter to ensure $C$ is ‘compact’ in e.g. the quaternion example). Now $L\subseteq C$ in general. It seems unlikely though in the quaternion case that $C$ coincides with the prepolar of the minimal nc convex set inside $L^{\circ}$, similarly to Effros and Winkler’s case.
7. Noncommutative Functions
7.1. Real NC Functions
Let $K$ be a real compact nc convex set. A nc function is a map $f:K\to\mathcal{M}(\mathbb{R})$ that is graded, preserves direct sums, and is unitarily equivariant. More specifically, it satisfies the following properties
  1. (1)
    $f(K_{n})\subseteq M_{n}(\mathbb{R})$
  2. (2)
    $f(\sum\alpha_{i}x_{i}\alpha^{\mathsf{T}}_{i})=\sum\alpha_{i}f(x_{i})\alpha^{\mathsf{T}}_{i}$ for every family of $\{x_{i}\in K_{n_{i}}\}$ and collection of isometries $\{\alpha_{i}\in M_{n,n_{i}}(\mathbb{R})\}$ such that $\sum\alpha_{i}\alpha_{i}^{\mathsf{T}}=1_{n}$.
  3. (3)
    $f(\beta x\beta^{\mathsf{T}})=\beta f(x)\beta^{\mathsf{T}}$ for every $x\in K_{n}$ and unitary $\beta\in M_{n}(\mathbb{R})$.
(Note that (3) is in fact a special case of (2).) We say that $f$ is bounded if it is uniformly bounded for all $k\in K$. The space of all real bounded nc functions on $K$ is $B(K)$. This has the uniform norm
$||f||=\sup_{k\in K}||f(k)||.$
We can similarly define $B(K,L)$ for $K,L$ real nc convex sets to be the nc functions from $K$ to $L$ which are bounded. Here, a nc function is the same definition as above, but with $\mathcal{M}(\mathbb{R})$ replaced by $L$.
As in [11], we have
Lemma 7.1.
If $K$ is a real compact nc convex set then $B(K)$ is a real $C^{*}$-algebra with the uniform norm and point-wise adjoint/product.
Proof.
The only difficulty in showing $B(K)$ is a $C^{*}$-algebra is showing it is complete. For the readers convenience we give the argument to show that it works in the real case. Let $f^{r}\in B(K)$ be a sequence such that $\sum_{r=1}^{\infty}||f^{r}||<\infty$. Define $f:K\to\mathcal{M}(\mathbb{F})$ by $f(x)=\sum_{r=1}^{\infty}f^{r}(x)$ for $x\in K_{m}$. This converges (absolutely) since $M_{n}(\mathbb{F})$ is complete. Because the $f^{r}$ are all graded and unitarily equivariant, $f$ will be too. Condition (3) in the definition above is easy to check. For (2), let $\{x_{i}\in K_{n_{i}}\}$ be set of elements in $K$, and $\alpha_{i}\in M_{n,n_{i}}(\mathbb{F})$ a family of isometries such that $\sum_{i}\alpha_{i}\alpha_{i}^{*}=1_{n}$. Take $\xi,\eta\in\ell_{n}^{2}$. Then, we have that
$\sum_{r=1}^{\infty}\sum_{i}|\langle\alpha_{i}f^{r}(x_{i})\alpha_{i}^{*}\xi,\eta\rangle|$ $\leq\sum_{n=1}^{\infty}\sum_{i}|\langle\alpha_{i}f^{r}(x_{i})\alpha_{i}^{*}\xi,\eta\rangle|$
$\leq\sum_{n=1}^{\infty}\big{(}\sum_{i}||f^{r}(x_{i})\alpha_{i}^{*}\xi)||^{2}\big{)}^{1/2}\big{(}\sum_{i}||\alpha_{i}^{*}\eta||^{2}\big{)}^{1/2}$
$\leq\sum_{n=1}^{\infty}||f^{r}||\big{(}\sum_{i}||\alpha_{i}^{*}\xi)||^{2}\big{)}^{1/2}\big{(}\sum_{i}||\alpha_{i}^{*}\eta||^{2}\big{)}^{1/2}$
$=\sum_{n=1}^{\infty}||f^{r}||\|\xi\|\|\eta\|,$
since e.g. $\big{(}\sum_{i}||\alpha_{i}^{*}\xi||^{2}\big{)}^{1/2}=\big{(}\sum_{i}\langle\alpha_{i}\alpha_{i}^{*}\xi,\xi\rangle\big{)}^{1/2}=||\xi||.$ Thus we may interchange the order of summation in $\sum_{r=1}^{\infty}\sum_{i}\,\langle\alpha_{i}f^{r}(x_{i})\alpha_{i}^{*}\xi,\eta\rangle$. We see that
$\langle f(\sum_{i}\,\alpha_{i}x_{i}\alpha^{*}_{i})\xi,\eta\rangle=\langle\sum\alpha_{i}f(x_{i})\alpha^{*}_{i}\xi,\eta\rangle$
as desired so that (2) holds. ∎
As in the complex case, $A(K)\hookrightarrow B(K)$ and we define $C(K)$ to be the $C^{*}$-algebra generated by $A(K)$ in $B(K)$.
7.2. Maximal C*-algebra
Let $S$ be a real (or complex) operator system. The maximal $C^{*}$-algebra generated by $S$, denoted $C^{*}_{\mathrm{max}}(S)$ is the $C^{*}$-algebra satisfying the following universal property:
${S}$${C^{*}_{\max}(S)}$${A}$$\scriptstyle{\iota}$$\scriptstyle{\varphi}$$\scriptstyle{\pi}$
where $\iota$ is a ucoi into a $C^{*}$-algebra $A$ such that $C^{*}(\iota(S))=C^{*}_{\mathrm{max}}(S)$, $\varphi$ is a ucoe such that $C^{*}(\varphi(S))=A$, and $\pi$ is an induced *-homomorphism.
Lemma 7.2.
[6, Lemma 5.1] For a real operator system $S$ we have $C^{*}_{\mathrm{max}}(S_{c})\cong C^{*}_{\mathrm{max}}(S)_{c}$ where the prior is the complex maximal $C^{*}$-algebra of the complex operator system $S_{c}$ and the latter is complexification of the real maximal $C^{*}$-algebra of $S$.
For any compact nc convex set $K$ we have $K\cong\mathrm{ncS}(A(K))$, and therefore any $k\in K_{n}$ corresponds to a nc state from $A(K)$ to $M_{n}$. The universal property of $C^{*}_{\text{max}}(A(K))$ gives a *-homomorphism $\delta_{x}:C^{*}_{\rm{max}}(A(K))\to M_{n}$ such that $\delta_{x}\circ\iota=\hat{x}$. Taking the double adjoint gives a normal *-homomorphism $\delta_{x}^{**}:C^{*}_{\rm{max}}(A(K))^{**}\to M_{n}$ and with this we define the map $\sigma:C^{*}_{\rm{max}}(A(K))^{**}\to B(K)$ by
$\sigma(b)(x)=\delta_{x}^{**}(b)$
for $b\in C^{*}_{\rm{max}}(A(K))^{**}$ and $x\in K$. Theorem $4.3.3$ of [11] shows that for a complex compact nc convex set, $B(K)$ is von-Neumann algebra isomorphic to $C^{*}_{\max,\mathbb{C}}(A(K))^{**}$ via the map $\sigma$ using the theory of Takesaki and Bichteler. Their proof also shows that $\sigma$ restricts to an isomorphism between $C^{*}_{\max,\mathbb{C}}(A(K))$ and $C(K)$ and that elements of $C(K)$ are the point-strong continuous nc functions on $K$.
Rather than using Takesaki-Bichteler theory, we can instead prove the real analogue of Theorem $4.3.3$ using complexification. Therefore, we need the following lemmas:
Lemma 7.3.
For real nc convex sets $K,L$, every real bounded nc map $f:K\to L$ has a unique complex bounded nc extension $f_{c}:K_{c}\to L_{c}$. If $L$ is complex bounded nc affine there is a complex bounded nc extension $K_{c}\to L$. These extensions are strongly continuous if $f$ is strongly continuous.
Proof.
For $x+iy\in(K_{c})_{n}$ define
$f_{c}(x+iy)=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,f(c(x+iy))\,\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
This function is a complex nc function by a proof similar to 3.8. First, it is clearly graded. Then, for $\beta\in M_{n}(\mathbb{C})$ a unitary and $x+iy\in(K_{c})_{n}$ we have
$f_{c}(\beta(x+iy)\beta^{*})$ $=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,f(c(\beta)c(x+iy)c(\beta)^{\mathsf{T}})\,\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,c(\beta)f(c(x+iy))c(\beta)^{\mathsf{T}}\,\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
$=\beta\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,f(c(x+iy))\,\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}\beta^{*}$
$=\beta f_{c}(x+iy)\beta^{*}$
which shows $(3)$. A similar proof shows $(2)$. In the case that $f$ is bounded, then $f_{c}$ will be bounded by Ruan’s first axiom. If $f$ is SOT continuous then $f_{c}$ will be too because the function $c$ is SOT continuous and so is adjoining the matrix $[1_{n};-i\cdot 1_{n}]$.
This will be the unique extension because for any complex nc function on $K_{c}$ extending $f$, say $g$, we have that
$(x+iy)\oplus(x-iy)=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\\ i\cdot 1_{n}&1_{n}\end{bmatrix}c(x+iy)\begin{bmatrix}1_{n}&-i\cdot 1_{n}\\ -i\cdot 1_{n}&1_{n}\end{bmatrix}$
so
$g(x+iy)\oplus g(x-iy)$ $=g\Big{(}\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\\ i\cdot 1_{n}&1_{n}\end{bmatrix}c(x+iy)\begin{bmatrix}1_{n}&-i\cdot 1_{n}\\ -i\cdot 1_{n}&1_{n}\end{bmatrix}\Big{)}$
$=\frac{1}{2}\begin{bmatrix}1_{n}&i\cdot 1_{n}\\ i\cdot 1_{n}&1_{n}\end{bmatrix}f(c(x+iy))\begin{bmatrix}1_{n}&-i\cdot 1_{n}\\ -i\cdot 1_{n}&1_{n}\end{bmatrix}.$
Here we first used that $g(x\oplus y)=g(x)\oplus g(y)$. The last part of the above equation is a $2n\times 2n$ matrix with top-right and bottom-left corners being 00. Comparing the top-left corners of the matrices in the above equation gives
$g(x+iy)=\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}&i\cdot 1_{n}\end{bmatrix}\,f(c(x+iy))\,\frac{1}{\sqrt{2}}\begin{bmatrix}1_{n}\\ -i\cdot 1_{n}\end{bmatrix}$
as desired. ∎
Lemma 7.4.
$B(K_{c})\cong B(K)_{c}$ as complex von Neumann algebras.
Proof.
We use the same maps as in Lemma 3.8, where we have a map
$\Gamma:B(K_{c})\to B(K)_{c}$
$\omega\mapsto\text{Re }\circ\omega_{|K}+i\,\text{Im }\circ\omega_{|K}$
with inverse $\Psi$. To see that $\Gamma$ is well-defined note
$||\text{Re }(\omega(x+i0))||\leq||\omega(x)||\leq||\omega||\qquad x\in K,$
and similarly for the imaginary part. Therefore,
$||c(\text{Re }\omega+i\text{Im }\omega)||\leq||\text{Re }\omega||+||\text{Im }\omega||\leq 2||\omega||$
A similar proof to that of Lemma 2.5 shows that ${\rm Re}$ and Im are unitarily equivariant. Also $\Gamma$ is a unital complete order isomorphism as may be shown similarly to the proof of Lemma 3.8, and therefore is a *-isomorphism. All *-isomorphisms are automatically normal and therefore $B(K_{c})\cong B(K)_{c}$ via $\Gamma$. Note that the proof that the inverse of $\Gamma$ is $\Psi$ does not work here because in showing $\Psi\circ\Gamma=\text{Id}$ we used compressions, however we can adjust the proof to use unitaries like at the end of 7.3 to make it work. ∎
Proposition 7.5.
$C(K_{c})\cong C(K)_{c}$ as complex $C^{*}$-algebras.
Proof.
We have
$C(K_{c})=C^{*}(A(K_{c}))\cong C^{*}(A(K)_{c})\cong C^{*}(A(K))_{c}=C(K)_{c}.$
We used the fact that if $A$ is a subsystem of a real $C^{*}$-algebra $B$ then $C^{*}(A_{c})=C^{*}(A)+iC^{*}(A)$ in $B_{c}$. ∎
Lemma 7.6.
Let $S(K)$ be the point-strong continuous functions in $B(K)$. Then $S(K_{c})\cong S(K)_{c}$ as $C^{*}$-algebras coming from the congruence $B(K_{c})\cong B(K)_{c}$
Proof.
We just need that the map $\Gamma$ in 7.4 satisfies $\Gamma(S(K_{c}))=S(K)_{c}$. Let $\omega\in S(K_{c})$ and $x_{\lambda}\xrightarrow[]{}x\in K_{n}$. Because Re is a contraction Re ($\text{Re }(\omega(x_{\lambda}+i0))\to\text{Re }(\omega(x+i0))$ in the strong operator topology. Similarly for Im. Therefore,
[…]
Thus $\Gamma(\omega)\in S(K)_{c}$. The converse is similar using the map $\Psi$ in Lemma 3.6. ∎
Theorem 7.7 (Real case of Theorem 4.3.3).
Let $K$ be a real compact nc convex set. Then the map $\sigma:C^{*}_{\rm{max}}(A(K))^{**}\to B(K)$ is a real linear normal *-isomorphism, which restricts to a *-isomorphism from $C^{*}_{\rm{max}}(A(K))$ onto $C(K)$. The elements of $C(K)$ are the point-strong continuous nc functions on $K$. Also, $\sigma\circ\iota$ is the identity map on $A(K)$.
Proof.
For $\mathcal{A}$ a real $C^{*}$-algebra, we have $(\mathcal{A}_{c})^{*}\cong(\mathcal{A}^{*})_{c}$ by [27]. The ensuing map $(\mathcal{A}_{c})^{**}\to(\mathcal{A}^{**})_{c}$ is a unital complete order isomorphism and normal *-isomorphism of complex von Neumann algebras [18]. Using this, Lemmas 7.2 and 3.8 give
$(C^{*}_{\rm{max},\mathbb{R}}(A(K))^{**})_{c}$ $\cong(C^{*}_{\rm{max},\mathbb{R}}(A(K))_{c})^{**}$
$\cong C^{*}_{\rm{max},\mathbb{C}}(A(K)_{c})^{**}$
$\cong C^{*}_{\rm{max},\mathbb{C}}(A(K_{c}))^{**}$
The complex case of this theorem and Lemma 7.4 gives
$C^{*}_{\rm{max},\mathbb{C}}(A(K_{c}))^{**}\cong B(K_{c})\cong B(K)_{c},$
with the composition $C^{*}_{\rm{max},\mathbb{C}}(A(K))^{**}\cong B(K)_{c}\cong B(K_{c})$ being via a (complex) normal *-isomorphism $\pi$ say. There is a real embedding of $C^{*}_{\rm{max},\mathbb{R}}(A(K))^{**}$ into $(C^{*}_{\rm{max},\mathbb{R}}(A(K))^{**})_{c}$, and similarly for $B(K)$ into $B(K)_{c}$. A diagram chase shows that the restriction of the complex normal *-isomorphism above is a real normal *-isomorphism $C^{*}_{\rm{max},\mathbb{R}}(A(K))^{**}\cong B(K)$, which is the ‘identity map’ on the copies of $A(K)$. From this we see that $B(K)$ is a von Neumann algebra.
To do the diagram chase, take $f\in A(K)$ which then embeds into $C^{*}_{\rm{max},\mathbb{R}}(A(K))^{**}$ and is denoted by $\theta(f)$. Going ’up’ in the diagram gives $\theta(f)+i0\in(C^{*}_{\rm{max},\mathbb{R}}(A(K))^{**})_{c}$. The isomorphism $(C^{*}_{\rm{max},\mathbb{R}}(A(K))^{**})_{c}\cong C^{*}_{\rm{max},\mathbb{C}}(A(K)_{c})^{**}$ will take $\theta(f)+i0$ to $\theta(f+i0)$. The third congruence in the above centered equations takes our element to $\theta(\psi(f+i0))$. Taking $\sigma$ of this element and evaluating at $x+i0\in K_{c}$ for $x\in K_{n}$ will give an element of $M_{n}(\mathbb{C})^{**}$. We evaluate this functional at $A\in M_{n}(\mathbb{C})^{*}$ and get
$\sigma(\theta(\psi(f+i0)))(x+i0)(A)$ $=\delta_{x+i0}^{**}(\theta(\psi(f+i0)))(A)$
$=A(\delta_{x+i0}(\psi(f+i0)))$
$=A(\psi(f+i0)(x+i0))$
$=A(f(x))$
On the other hand,
$\sigma(\theta(f))(x)(A)=\delta_{x}^{**}(\theta(f))(A)=A(f(x))$
Therefore, the diagram chase shows that it is the ‘identity map’ on the copies of $A(K)$ that extends to $\pi$. Since $\pi(A(K))$ is the copy of $A(K)$ in $B(K)$ inside $B(K_{c})$, it follows that $\pi(C^{*}_{\rm{max}}(A(K))^{**})$ is a $C^{*}$-subalgebra $D$ of $B(K)$ with $D+iD=B(K)_{c}$. Hence $D=B(K)=\pi(C^{*}_{\rm{max}}(A(K))^{**})$. Also $I_{A(K)}$ extends to a *-isomorphism between the $C^{*}$-algebra generated by $A(K)$ in both sets, so $C^{*}_{\rm{max}}(A(K))\cong C(K)$. It also shows that $\sigma\circ\iota$ is the identity map on $A(K)$. That $C(K)$ are the point-strong continuous nc functions follows from the complex case and 7.6. Indeed $C(K)=B(K)\cap C(K)_{c}=B(K)\cap C(K_{c})$. ∎
As we saw after Lemma 7.2, any element $k$ of $K$ defines a nc/matrix state on $C(K)$, and a weak* continuous matrix state on $B(K)$. In particular, $f\mapsto f(k)$ is weak* continuous on $B(K)$.
Corollary 7.8.
Let $K$ be a real compact nc convex set. The real enveloping von Neumann algebra $C(K)^{**}$ of $C(K)$ is isomorphic to the real $C^{*}$-algebra $B(K)$ of bounded nc functions on $K$. The real dual operator system $A(K)^{**}$ is completely order isomorphic to the real operator system of bounded nc affine functions on $K$. The latter space has as complexification the bounded complex nc affine functions on $A_{c}$.
Proof.
To see that $A(K)^{**}$ is completely order isomorphic to the real operator system of bounded nc affine functions on $K$, note that $A(K)^{\perp\perp}\subseteq B(K)$. We need to show $f\in B(K)$ is in $A(K)^{\perp\perp}$ if and only if $f(\beta^{*}x\beta)=\beta^{*}f(x)\beta$ for every $x\in K_{m}$ and isometry $\beta\in M_{m,n}(\mathbb{R})$. Since this the latter is true for $f\in A(K)$, it will also be true by a weak* approximation argument for $f\in A(K)^{\perp\perp}$, using the fact above the corollary. This gives a weak* continuous ucoe $\nu:A(K)^{\perp\perp}\to BA(K)$, where $BA(K)$ is the operator subsystem of bounded nc affine functions in $B(K)$. Then $BA(K)_{c}=BA(K)+iBA(K)=BA_{\mathbb{C}}(K_{c})$ in $B(K)_{c}=B(K_{c})$. The rest follows by complexification from Corollary 4.3.3 in [11]. Indeed if $\nu$ were not surjective then its complexification would also not be, contradicting that $BA_{\mathbb{C}}(K_{c})\cong A_{\mathbb{C}}(K_{c})^{\perp\perp}$. ∎
The proof of the following is the same as in the complex case (see Proposition 2.5.3 of [11]).
Proposition 7.9.
Let $K$ be a real compact nc convex set and $f\in C(K)$ a continuous nc function. Then $f$ is bounded with
$||f||=\sup_{n<\infty}||f|_{K_{n}}||.$
7.3. Minimal $C^{*}$-algebra
As in the complex case, every real operator system $V$ has a $C^{*}$-envelope or minimal $C^{*}$-algebra denoted by $C^{*}_{\min}(V)$. There is a ucoe $\iota:V\to C^{*}_{\min}(V)$ which satisfies the following universal property:
${A=C^{*}(\varphi(V))}$${V}$${C^{*}_{\min}(V)}$$\scriptstyle{\pi}$$\scriptstyle{\varphi}$$\scriptstyle{\iota}$
where $\varphi$ is a ucoe of $V$ into another real $C^{*}$-algebra $A$ such that $\varphi(V)$ generates $A$ as a $C^{*}$-algebra, and $\pi$ is an induced surjective *-homomomorphism making the diagram commute.
Lemma 7.10.
[4, Corollary 4.3] For a real operator system $S$ we have $C^{*}_{\min}(S_{c})\cong C^{*}_{\min}(S)_{c}$ where the prior is the complex minimal $C^{*}$-algebra of the complex operator system $S_{c}$ and the latter is complexification of the real minimal $C^{*}$-algebra of $S$.
Example 7.11.
If $\mathcal{A}$ is a real/complex $C^{*}$-algebra viewed as an operator system, the universal property shows $C^{*}_{\min}(\mathcal{A})=\mathcal{A}$. There exists $C^{*}$-algebras that are not the complexification of real $C^{*}$-algebras by [25, Problem 1.5] and we can use this to construct complex operator systems and compact nc convex sets which are not complexifications. For instance, let $A$ be such a complex $C^{*}$-algebra viewed as an operator system, then if it were the complexification of a real operator system $V$ we would get
$\mathcal{A}=C^{*}_{\min}(\mathcal{A})=C^{*}_{\min}(V_{c})\cong C^{*}_{\min}(V)_{c}$
and the latter is the complexification of a real $C^{*}$-algebra which is a contradiction. Similarly, if every complex compact nc convex set was a complexification, then $\mathrm{ncS}(\mathcal{A})$ would be $L_{c}$ for some real compact nc convex set $L$. By Theorem 4.1 we have that $\mathcal{A}\cong A(L_{c})$ as complex operator systems and so
$\mathcal{A}\cong C^{*}_{\min}(A(L_{c}))\cong C^{*}_{\min}(A(L))_{c},$
which is a contradiction.
Acknowledgements
We acknowledge support from NSF Grant DMS-2154903. We thank Travis Russell for several conversations and helpful thoughts and information, and thank Matt Kennedy who very very recently, in answer to a couple of questions of ours related to our draft, kindly drew our attention to some items in the ArXiV revision of [11] of June 10, 2025.